### v1

# Global well-posedness of non-integrable hyperbolic-ellptic Ishimori system in the critical Sobolev space

Abstract.
We consider the Cauchy problem for the hyperbolic-elliptic Ishimori system with general decoupling constant $\kappa\in\mathbb{R}$ and prove global well-posedness in the critical Sobolev space. The proof relies primarily on new bilinear estimates, which are established via a novel div-curl lemma first introduced by the second author in [14]. Our approach combines the caloric gauge technique with $U^{p}$-$V^{p}$ type Strichartz estimates to handle the hyperbolic structure of the equation. The results extend previous work on the integrable case $\kappa=1$ to general $\kappa$ and provide a unified framework that also applies to hyperbolic and elliptic Schrödinger maps in dimensions $d\geq 2$.
1. Introduction
We consider the hyperbolic-elliptic Ishimori system, a two-dimensional topological spin field model with the form
(1.1) $\begin{cases}\partial_{t}S=S\times(\partial_{x_{1}}^{2}S-\partial_{x_{2}}^{2}S)+\kappa(\partial_{x_{1}}\phi\cdot\partial_{x_{1}}S+\partial_{x_{2}}\phi\cdot\partial_{x_{2}}S),&\text{on }\mathbb{R}^{2}\times\mathbb{R},\\ \Delta_{x}\phi=2S\cdot(\partial_{x_{1}}S\times\partial_{x_{2}}S),\\ S|_{t=0}=S_{0},\end{cases}$
where $S:\mathbb{R}^{2}\times\mathbb{R}\to\mathbb{S}^{2}$ is a spin field taking values in the unit sphere, $\phi:\mathbb{R}^{2}\times\mathbb{R}\to\mathbb{R}$ is a scalar potential, and $\kappa\in\mathbb{R}$ is the coupling constant. The system was introduced by Ishimori in [5] as a two-dimensional generalization of the two dimensional Heisenberg equation in ferromagnetism. The potential $\phi$ is related to the topological charge density $2S\cdot(\partial_{x_{1}}S\times\partial_{x_{2}}S)$. The total topological charge, defined by
(1.2) $Q=\frac{1}{4\pi}\int_{\mathbb{R}^{2}}S\cdot(\partial_{x_{1}}S\times\partial_{x_{2}}S),$
represents the degree of the mapping $S:\mathscr{T}(\cong\mathbb{R}^{2}\cup\{\infty\})\to\mathscr{T}$.
The system exhibits different behaviors depending on the coupling constant. When $\kappa=0$, the system reduces to the two-dimensional hyperbolic Heisenberg equation (hyperbolic Schrödinger map). When $\kappa=1$, the system is completely integrable and can be solved via the inverse scattering method. For general $\kappa\in\mathbb{R}$, the system is non-integrable and presents additional analytical challenges.
The local and global regularity properties of the Cauchy problem associated with the hyperbolic-elliptic Ishimori system have been extensively studied, see [6, 9, 13]. Soyeur [9] established local and global existence for small initial data in $H^{3}$, and uniqueness for large data in $H^{4}$. Wang [13] proved local well-posedness for small data in $H^{\sigma}_{Q}$ for $\sigma>\frac{3}{2}$. For the integrable case $\kappa=1$, Bejenaru, Ionescu, and Kenig [6] established global well-posedness for small data in the critical Sobolev space $\dot{H}^{1}_{Q}(\mathbb{R}^{2})$. For a more comprehensive historical overview, we refer to [8, Section 9.2].
Regarding the elliptic Schrödinger map equation, which relates to the special case $\kappa=0$ in (1.1), substantial progress has been made on low-regularity local and global well-posedness, see [1, 3, 4]. Of particular relevance to our approach is the seminal work of Bejenaru, Ionescu, Kenig, and Tataru [1], who proved global existence for the elliptic Schrödinger map in critical Sobolev spaces for dimensions $d\geq 2$. Their proof relies crucially on the caloric gauge formulation and local smoothing estimates for the associated linear flow.
This paper aims to extend these results to the non-integrable Ishimori system ($\kappa\in\mathbb{R}$) using the caloric gauge framework from [1]. However, there are significant differences: the local smoothing estimates used in the elliptic case are not directly applicable due to the hyperbolic structure of the Ishimori equation. Instead, we employ $U^{p}$-$V^{p}$ type Strichartz estimates and a new div-curl lemma first introduced by the second author in [14]. This novel div-curl lemma is crucial for establishing the necessary bilinear estimates, particularly for controlling high-low and low-high frequency interactions in Bony’s paraproduct decomposition. We note that our methods are also applicable to both hyperbolic and elliptic Schrödinger maps in dimensions $d\geq 2$, providing a unified approach.
Before stating our main results, we introduce some notations. For $\sigma\in[0,\infty)$, let $H^{\sigma}$ denote the usual Sobolev space of complex valued function. Given a point $Q\in\mathbb{S}^{2}$, we define the Sobolev space $H^{\sigma}_{Q}$ by
(1.3) $H^{\sigma}_{Q}:=\{f:\mathbb{R}^{2}\to\mathbb{R}^{3}\ |\ |f|\equiv 1\ \text{a.e. and}\ f-Q\in H^{\sigma}\},$
which is equipped with the metric $d_{H^{\sigma}_{Q}}(f,g)=\|f-g\|_{H^{\sigma}}$. Similarly for $\dot{H}^{\sigma}_{Q}$ and $H^{\sigma,\rho}_{Q}$. We define
(1.4) $H_{Q}^{\infty}=\bigcap_{k=1}^{\infty}H^{k}_{Q}.$
Our main result in this paper is the following small data global well-posedness.
Theorem 1.1.
Given $Q\in\mathbb{S}^{2}$. Then there exists $\varepsilon>0$ such that for any $S_{0}\in H^{\infty}_{Q}$ with $\|S_{0}-Q\|_{\dot{H}^{1}}\leq\varepsilon$, there exists a unique global solution $S\in C(\mathbb{R};H^{\infty}_{Q})$ to the ishimori system (1.1) satisfying
(1.5) $\sup_{t}\|S(t)-Q\|_{\dot{H}^{1}}\lesssim\|S_{0}-Q\|_{\dot{H}^{1}},$
and for $k\in\mathbb{Z}_{+},$
(1.6) $\sup_{t}\|S(t)\|_{H_{Q}^{k}}\leq C(k,\|S_{0}\|_{H^{k}_{Q}}).$
Moreover, for any $\sigma\in[0,\sigma_{1}]$ the operator $T_{Q}:S_{0}\mapsto S(t)$ admits a continuous extension
(1.7) $T_{Q}:\mathcal{B}^{\sigma}_{\varepsilon(\sigma_{1})}\to C(\mathbb{R},H_{Q}^{\sigma+1}),$
where
(1.8) $\mathcal{B}^{\sigma}_{\varepsilon(\sigma_{1})}:=\{f\in H^{\sigma+1}_{Q}\ |\ \|f-Q\|_{\dot{H}^{1}}\leq\varepsilon(\sigma_{1})\}.$
1.1. The modified spin model in caloric gauge
Following the procedure in [1], we construct the fields $\psi_{m}$ and the connection coefficients $A_{m}$, and derive the differentiated Ishimori equation satisfied by these functions. To fix the connection coefficients uniquely, we choose the caloric gauge, which is implemented by solving a heat equation and thus extending the spin field $S$ to include an auxiliary parabolic time variable $s\in[0,\infty)$.
Instead of working directly on the spin field $S$, we study its derivatives $\partial_{\alpha}S$ for $\alpha=t,1,2$, which are tangent vectors in $T_{S(x,t)}\mathbb{S}^{2}$. Now suppose that there exists a smooth frame $(v,w)=(v,S\times v)\in T_{S(x,t)}\mathbb{S}^{2}$. We introduce the complexified differentiated variables
(1.9) $\psi_{\alpha}=v\cdot\partial_{\alpha}S+iw\cdot\partial_{\alpha}S,$
and the real connection coefficients
(1.10) $A_{\alpha}=w\cdot\partial_{\alpha}v.$
Since the vectors $(S,v,w)$ form a orthonormal frame for $T\mathbb{R}^{3}$, it follows that
(1.11) $\begin{cases}\partial_{\alpha}S=v\Re(\psi_{\alpha})+w\Im(\psi_{\alpha}),\\ \partial_{\alpha}v=-S\Re(\psi_{\alpha})+wA_{\alpha},\\ \partial_{\alpha}w=-S\Im(\psi_{\alpha})-vA_{\alpha}.\end{cases}$
Using the above formulas, one can verify that $\psi_{\alpha}$ and $A_{\alpha}$ satisfy the covariant curl relations
(1.12) $(\partial_{\alpha}+iA_{\alpha})\psi_{\beta}=(\partial_{\beta}+iA_{\beta})\psi_{\alpha}.$
Introducing the covariant derivative $D_{\alpha}=\partial_{\alpha}+iA_{\alpha}$, we can rewrite this as
(1.13) $D_{\alpha}\psi_{\beta}=D_{\beta}\psi_{\alpha}.$
Direct computation shows that
(1.14) $\partial_{\alpha}A_{\beta}-\partial_{\beta}A_{\alpha}=\Im(\psi_{\alpha}\overline{\psi_{\beta}})=:q_{\alpha\beta},$
(1.15) $[D_{\alpha},D_{\beta}]=iq_{\alpha\beta},$
where $q_{\alpha\beta}$ represents the curvature of the connection.
We now express the original Ishimori system (1.1) in terms of $\psi_{\alpha}$ and $A_{\alpha}$.
For the Poisson equation in (1.1), we compute using (1.11)
$\Delta_{x}\phi$ $=2S\cdot[(v\Re(\psi_{1})+w\Im(\psi_{1}))\times(v\Re(\psi_{2})+w\Im(\psi_{2}))]$
$=2S\cdot(v\times w)(\Re(\psi_{1})\Im(\psi_{2})-\Re(\psi_{2})\Im(\psi_{1}))$
$=2\Im(\psi_{2}\overline{\psi_{1}})=2(\partial_{2}A_{1}-\partial_{1}A_{2})=-2\epsilon_{ij}\partial_{i}A_{j}.$
Here and in what follows we use the notation $\epsilon_{ij}=\delta_{1i}\delta_{2j}-\delta_{1j}\delta_{2i}$ and the summation convention over repeated indices. Hence,
(1.16) $\phi=(-\Delta)^{-\frac{1}{2}}\epsilon_{ij}R_{i}A_{j},\qquad\partial_{m}\phi=\epsilon_{ij}R_{m}R_{i}A_{j}$
with $R_{i}=(-\Delta)^{-\frac{1}{2}}\partial_{i}$ denoting the Riesz transforms.
For the evolution equation, substituting (1.11) into (1.1) gives
$\partial_{t}S$ $=S\times[v(\partial_{1}\Re\psi_{1}-A_{1}\Im\psi_{1})+w(\partial_{1}\Im\psi_{1}+A_{1}\Re\psi_{1})]$
$\ \ \ -S\times[v(\partial_{2}\Re\psi_{2}-A_{2}\Im\psi_{2})+w(\partial_{2}\Im\psi_{2}+A_{2}\Re\psi_{2})]$
$\ \ \ +\kappa[v\Re(\psi_{1})+w\Im(\psi_{1})]\partial_{1}\phi+\kappa[v\Re(\psi_{2})+w\Im(\psi_{2})]\partial_{2}\phi$
$=v(-\partial_{1}\Im\psi_{1}-A_{1}\Re\psi_{1}+\partial_{2}\Im\psi_{2}+A_{2}\Re\psi_{2}+\kappa\Re(\psi_{1})\partial_{1}\phi+\kappa\Re(\psi_{2})\partial_{2}\phi)$
$\ \ \ +w(\partial_{1}\Re\psi_{1}-A_{1}\Im\psi_{1}-\partial_{2}\Re\psi_{2}+A_{2}\Im\psi_{2}+\kappa\Im(\psi_{1})\partial_{1}\phi+\kappa\Im(\psi_{2})\partial_{2}\phi).$
Using $\psi_{t}=v\cdot\partial_{t}S+iw\cdot\partial_{t}S$, we obtain the compact expression
(1.17) $\psi_{t}=v\cdot\partial_{t}S+iw\cdot\partial_{t}S=i(D_{1}\psi_{1}-D_{2}\psi_{2})+\kappa\psi_{l}\partial_{l}\phi,$
which expresses the time derivative of the spin field in terms of the spatial derivatives and connection coefficients.
Applying the compatibility condition (1.12), we derive the evolution equations for $\psi_{m}$ ($m=1,2$) as
$iD_{t}\psi_{m}$ $=iD_{m}\psi_{t}=-D_{m}(D_{1}\psi_{1}-D_{2}\psi_{2})+i\kappa D_{m}(\psi_{l}\partial_{l}\phi)$
$=-D_{m}(D_{1}\psi_{1}-D_{2}\psi_{2})+i\kappa(D_{m}\psi_{l}\partial_{l}\phi+\psi_{l}\partial_{l}\partial_{m}\phi)$
(1.18) $=-(D_{1}^{2}-D_{2}^{2})\psi_{m}-i(q_{m1}\psi_{1}-q_{m2}\psi_{2})+i\epsilon_{ij}\kappa(D_{m}\psi_{l}R_{l}R_{i}A_{j}+\psi_{l}\partial_{m}(R_{l}R_{i}A_{j})).$
Expanding the covariant derivatives yields the following nonlinear ultrahyperbolic Schrödinger equation
(1.19) $(i\partial_{t}+\mu_{l}\partial_{l}^{2})\psi_{m}=$ $-2i\mu_{l}A_{l}\partial_{l}\psi_{m}+(A_{t}+\mu_{l}(A_{l}^{2}-i\partial_{l}A_{l}))\psi_{m}-i\mu_{l}\psi_{l}\Im(\psi_{m}\bar{\psi}_{l})$
$+i\epsilon_{ij}\kappa(D_{m}\psi_{l}R_{l}R_{i}A_{j}+\psi_{l}\partial_{m}(R_{l}R_{i}A_{j})),$
where $\mu_{l}=\delta_{1l}-\delta_{2l}$ reflects the ultrahyperbolic signature.
Thus, The system (1.12), (1.14), (1.17), and (1.19) constitutes the modified spin model, a formulation of the Ishimori system in terms of the gauge-dependent variables $\psi_{\alpha}$ and $A_{\alpha}$.
To obtain a well-posed system, we impose Tao’s caloric gauge condition, defined as follows:
Definition 1.2 (Caloric gauge).
Let $S$ be a solution of (1.1) in $C(\mathbb{R};H^{\infty}_{Q})$ and $(v_{\infty},w_{\infty})=(v_{\infty},Q\times v_{\infty})$ be the orthonormal frame in $T_{Q}\mathbb{S}^{2}$. A caloric gauge is a tuple consisting of a extended map $\widetilde{S}:\mathbb{R}_{+}\times\mathbb{R}\times\mathbb{R}^{2}\to\mathbb{S}^{2}$ and an orthonormal frame $(v,w)=(v,\widetilde{S}\times v)$ for $T_{\widetilde{S}}\mathbb{S}^{2}$ such that
(1.20) $\begin{cases}\partial_{s}\widetilde{S}=\Delta_{x}\widetilde{S}+\widetilde{S}\sum_{m=1,2}|\partial_{m}\widetilde{S}|^{2},\ \text{for}\ s\in[0,\infty)\\ \widetilde{S}(0,t,x)=S(t,x),\end{cases}$
and the following gauge conditions hold:
(1.21) $A_{s}:=w\cdot\partial_{s}v=0,\ \lim_{s\to\infty}(v,w)=(v_{\infty},w_{\infty}).$
The existence and uniqueness of such a gauge for small initial data are guaranteed by Lemma 1.5.
We now define the extended variables $\psi_{\alpha^{\prime}},A_{\alpha^{\prime}},\ \alpha^{\prime}=s,t,1,2$ by
(1.22) $\begin{cases}\psi_{\alpha^{\prime}}=v\cdot\partial_{\alpha^{\prime}}\widetilde{S}+iw\cdot\partial_{\alpha^{\prime}}\widetilde{S},\\ A_{\alpha^{\prime}}=w\cdot\partial_{\alpha^{\prime}}w\end{cases}$
The parallel transport condition in the gauge definition implies the key gauge condition
(1.23) $A_{s}=0.$
A similar computation shows that (1.20) is equivalent to
(1.24) $\psi_{s}=D_{1}\psi_{1}+D_{2}\psi_{2}.$
Taking $D_{\alpha}$ derivative and using (1.12) again, the heat equations for extended variables $\psi_{\alpha},\ \alpha=t,1,2,$ take the form of
(1.25) $(\partial_{s}-\Delta_{x})\psi_{\alpha}=2iA_{l}\partial_{l}\psi_{\alpha}-(A_{l}^{2}-i\partial_{l}A_{l})\psi_{\alpha}+i\Im(\psi_{\alpha}\overline{\psi_{l}})\psi_{l}.$
Moreover, from $A_{s}=0$ and (1.14) we have
(1.26) $\partial_{s}A_{\alpha}=\Im(\psi_{s}\overline{\psi_{\alpha}})=\Im(\overline{\psi_{\alpha}}D_{l}\psi_{l}),$
which, together with decay estimates (1.42) as $s\to\infty$, yields the integral representation
(1.27) $A_{\alpha}(s)=-\int_{s}^{+\infty}\Im(\overline{\psi_{\alpha}}D_{l}\psi_{l})(r)\mathrm{d}r,\ \alpha=t,1,2.$
To establish continuous dependence on initial data, we also require the linearized Ishimori equation. Consider a one-parameter family of solutions $S^{h}$ to (1.1) with $S^{h_{0}}=S$, and define $S_{\text{lin}}=\partial_{h}S^{h}|_{h_{0}}$. Linearizing (1.1) gives
(1.28) $\partial_{t}S_{\text{lin}}=S_{\text{lin}}\times\mu_{l}\partial^{2}_{l}S+S\times\mu_{l}\partial^{2}_{l}S_{\text{lin}}+\kappa\partial_{x}\phi_{\text{lin}}\cdot\partial_{x}S+\kappa\partial_{x}\phi\cdot\partial_{x}S_{\text{lin}},$
with the constraint $S_{\text{lin}}\cdot S=0$ (since variations preserve the sphere constraint). The linearized potential $\phi_{\text{lin}}$ satisfies
(1.29) $\Delta\phi_{\text{lin}}$ $=2[S_{\text{lin}}\cdot(\partial_{1}S\times\partial_{2}S)+S\cdot(\partial_{1}S_{\text{lin}}\times\partial_{2}S)+S\cdot(\partial_{{}_{1}}S\times\partial_{2}S_{\text{lin}})]$
$=2S\cdot[\partial_{1}(S_{\text{lin}}\times\partial_{2}S)-\partial_{2}(S_{\text{lin}}\times\partial_{1}S)]$
$=2\partial_{1}[S\cdot(S_{\text{lin}}\times\partial_{2}S)]-2\partial_{2}[S\cdot(S_{\text{lin}}\times\partial_{1}S)].$
Decomposing $S_{\text{lin}}$ in the frame as
(1.30) $S_{\text{lin}}=v\Re(\psi_{lin})+w\Im(\psi_{lin}),$
and repeating the earlier computations yields the linearized equation for $\psi_{\text{lin}}$
(1.31) $(i\partial_{t}+\mu_{l}\partial_{l}^{2})\psi_{\text{lin}}=$ $-2i\mu_{l}A_{l}\partial_{l}\psi_{\text{lin}}+(A_{t}+\mu_{l}(A_{l}^{2}-i\partial_{l}A_{l}))\psi_{\text{lin}}-i\mu_{l}\psi_{l}\Im(\psi_{\text{lin}}\bar{\psi}_{l})$
$~+i\kappa\epsilon_{ij}R_{l}R_{i}A_{j}\cdot D_{l}\psi_{\text{lin}}-i\kappa\psi_{l}R_{l}[R_{1}\Im(\overline{\psi_{2}}\psi_{\text{lin}})-R_{2}\Im(\overline{\psi_{1}}\psi_{\text{lin}})].$
1.2. Outline of the proof
We adopt the analytical framework introduced in [1] to establish our main results. Given a solution $S\in C(\mathbb{R};H^{\infty}_{Q})$ to the Ishimori system, our main goal is to prove a priori bound on $\|S\|_{L^{\infty}_{t}(\dot{H}^{1}\cap H^{\sigma+1}_{Q})}$ for $\sigma$ in a fixed interval $0,\sigma_{1}]$. We shall use the homogenous Littlewood-Paley decomposition and the notation of frequency envelopes.
Definition 1.3.
For $k\in\mathbb{Z}$, we define the standard homogenous Littlewood-Paley operator $P_{k}$. Let $\chi$ be the smooth cutoff to the region [1,1][-1,1], and $P_{k},P_{\leq k}$ are defined by
(1.32) $\widehat{P_{\leq k}f}(\xi):=\chi(2^{-k}|\xi|)\hat{f}(\xi),\ P_{k}=P_{\leq k}-P_{\leq\frac{k}{2}}.$
Definition 1.4.
A positive squence $\{b_{k}\}$ is a frequency envelope if is $\ell^{2}$ bounded
(1.33) $\sum_{k\in\mathbb{Z}}b_{k}^{2}<\infty,$
and slowly varying,
(1.34) $b_{k}\leq 2^{\delta|k-j|}b_{j},\ k,j\in\mathbb{Z},$
where $\delta$ is a sufficiently small positive parameter.
An $\epsilon$-frequency envelope satisfies additional condition
(1.35) $\sum_{k}b_{k}^{2}\leq\epsilon^{2}$
Given a $\ell^{2}$ positive sequence $\alpha_{k}$, we often define its related frequency envelope
(1.36) $\alpha^{\prime}_{k}=\sup_{j}2^{-\delta|j-k|}\alpha_{j}.$
It’s clear that $\alpha^{\prime}_{k}$ is indeed a frequency envelope satisfying
(1.37) $\alpha_{k}\leq\alpha^{\prime}_{k},\ \sum_{k}(\alpha^{\prime}_{k})^{2}\lesssim\sum_{k}\alpha_{k}^{2}.$
Let $S(s):=S(s,x,t)$ be the solution to the heat flow in the caloric gauge with initial data $S=S(x,t)\in C(\mathbb{R},H^{\infty}_{Q})$, for $\sigma\geq 1$ we introduce the frequency envelope related to $S$:
(1.38) $\gamma_{k}(\sigma)=\sup_{j\in\mathbb{Z}}2^{-\delta|k-j|}2^{(\sigma+1)j}\|P_{j}S(0)\|_{L_{t}^{\infty}L^{2}_{x}},\ \sigma\in[0,\sigma_{1}+1].$
we also let $\gamma_{k}:=\gamma_{k}(0)$. The existence of caloric gauge is ensured by the following lemma:
Lemma 1.5.
([1, Prop. 4.2])Given arbitary interval $I\subseteq\mathbb{R}$ and $S\in C(I;H^{\infty}_{Q})$ satisfying the smallness condition
(1.39) $\sum_{k\in\mathbb{Z}}2^{2k}\|P_{j}S\|^{2}_{L^{\infty}_{t}L^{2}_{x}}=\sum_{k\in\mathbb{Z}}\gamma_{k}^{2}\ll 1,$
then there exists a unique corresponding caloric gauge as defined in Definition 1.2. Moreover, we have the bounds
(1.40) $\|P_{k}(\widetilde{S},v,w)(s)\|_{L^{\infty}_{t}L^{2}_{x}}\lesssim\gamma_{k}(\sigma)\langle 2^{2k}s\rangle^{-20}2^{-(\sigma+1)k},\ \sigma\in[0,\sigma_{1}],$
and for $\sigma\in\mathbb{Z}_{+}$,
(1.41) $\sup_{k\in\mathbb{Z}}\sup_{s\in[0,\infty)}\langle s\rangle^{\frac{\sigma}{2}}2^{\sigma k}\|P_{k}(\widetilde{S},v,w)(s)\|_{L^{\infty}_{t}L^{2}_{x}}<\infty.$
As a conclusion we have
(1.42) $\sup_{k\in\mathbb{Z}}\sup_{s\in[0,\infty)}\langle s\rangle^{\frac{\sigma}{2}}2^{(\sigma-1)k}\|P_{k}(\psi_{m}(s),A_{m}(s))\|_{L^{\infty}_{t}L^{2}_{x}}<\infty,\ \text{for}\ m=1,2,$
$\sup_{k\in\mathbb{Z}}\sup_{s\in[0,\infty)}\langle s\rangle^{\frac{\sigma}{2}}2^{\sigma k}\|P_{k}(\psi_{t}(s),A_{t}(s))\|_{L^{\infty}_{t}L^{2}_{x}}<\infty.$
We now state our main bootstrap arguments. Given arbitary interval $I\subseteq\mathbb{R}$ and solution $S\in C(I;H^{\infty}_{Q})$ satisfying the smallness condition in Lemma 1.5, we shall work with the caloric gauge and the associated fields and connection coefficients $\psi_{\alpha},A_{\alpha}$.
For simplicity, we adopt the notation
(1.43) $\bm{\psi}=(\psi_{m},\overline{\psi_{m}})_{m=1,2},\ \bm{A}=(A_{m})_{m=1,2}.$
Denote by $G=L^{\infty}_{t}L^{2}_{x}\cap L^{4}_{t,x}$ the Strichartz space. In view of the bilinear estimate in Lemma 2.25, we introduce the functional
(1.44) $\mathcal{D}(u):=\|u\|_{G}+\sup_{y}\|u^{y}\cdot(i\partial_{t}+\mu_{l}\partial_{l}^{2})u\|_{L^{1}_{t,x}}.$
We now introduce three families of frequency envelopes that will govern our bootstrap analysis. For $\sigma\in[0,\sigma_{1}]$ and $m=1,2$ we define
(1.45) $a_{k}(\sigma)=\sup_{j\in\mathbb{Z}}2^{-\delta|k-j|}\sup_{s\geq 0}\langle 2^{2j}s\rangle^{4}(2^{\sigma j}\|P_{j}\bm{\psi}(s)\|_{G}+1_{\{\sigma\geq\frac{1}{5}\}}2^{(\sigma-1)j}\|P_{j}\psi_{t}(s)\|_{G}),$
(1.46) $b_{k}(\sigma)=\sup_{j\in\mathbb{Z}}2^{-\delta|k-j|}(2^{\sigma j}\|P_{j}\bm{\psi}(0)\|_{G}+1_{\{\sigma\geq\frac{1}{5}\}}2^{(\sigma-1)j}\|P_{j}\psi_{t}(0)\|_{G}),$
(1.47) $c_{k}(\sigma)=\sup_{j\in\mathbb{Z}}2^{-\delta|k-j|}2^{\sigma j}\|P_{j}\nabla S_{0}\|_{L^{2}_{x}}.$
Clearly,
(1.48) $c_{k}(\sigma)\leq b_{k}(\sigma)\leq a_{k}(\sigma).$
We use the notation
(1.49) $(a_{k},b_{k},c_{k}):=(a_{k}(0),b_{k}(0),c_{k}(0))$
to measure the critical regularity. Under the hypotheses of Theorem 1.1, $c_{k}$ forms an $\varepsilon$-frequency envelope.
Our proof strategy revolves around two key bootstrap propositions. The first controls the heat flow evolution.
Proposition 1.6 (Heat flow bootstrap assumptions).
Let $a_{k}(\sigma),b_{k}(\sigma)$ be defined as in (1.45) and (1.46). Suppose that $b_{k}$ is a $\varepsilon^{\frac{3}{4}}$-frequency envelope and
(1.50) $a_{k}(\sigma)\leq\varepsilon^{-\frac{1}{4}}b_{k}(\sigma),\ \mathcal{D}(P_{k}\bm{\psi})\leq 2^{-\sigma k}b_{k}(\sigma),\ \sigma\in[0,\sigma_{1}].$
Then we have the improved bounds
(1.51) $a_{k}(\sigma)$ $\lesssim b_{k}(\sigma),\ \sigma\in[0,\sigma_{1}]$
(1.52) $\sup_{y}\|P_{k_{1}}\bm{\psi}(s)P_{k_{2}}w^{y}\|_{L^{2}_{t,x}}$ $\lesssim 2^{-\frac{|k_{1}-k_{2}|}{2}}2^{-\sigma k}\langle 2^{2k}s\rangle^{-3}b_{k}(\sigma)\mathcal{D}(P_{k_{2}}w).$
The second bootstrap proposition concerns the ultrahyperbolic Schrödinger evolution.
Proposition 1.7 (Ultrahyperbolic Schrödinger bootstrap assumptions).
Let $b_{k}(\sigma),c_{k}(\sigma)$ be defined as in (1.46) and (1.47). Suppose that
(1.53) $\mathcal{D}(P_{k}\bm{\psi})\leq b_{k}(\sigma)\leq\varepsilon^{-\frac{1}{4}}c_{k}(\sigma),\ \sigma\in[0,\sigma_{1}].$
Then we have the improved bounds
(1.54) $b_{k}(\sigma)+\mathcal{D}(P_{k}\bm{\psi})\lesssim c_{k}(\sigma),\ \sigma\in[0,\sigma_{1}].$
Proposition 1.6 is proved in Section 3, while Proposition 1.7 is proved in Section 4. Following the strategy in [1], we show that Proposition 1.7 implies Theorem 1.1.
Proof of Theorem 1.1.
Our starting point is the local local-in-time existence and uniqueness of ishimori equation given by Kenig and Nahmod [6]: if $S_{0}\in H_{Q}^{\infty}$, then there exists $T=T(\|S_{0}\|_{H^{2}_{Q}})$ and a unique solution $S\in C((-T,T),H^{\infty}_{Q})$ of the Cauchy problem (1.1). Our goal is using Proposition 1.7 to prove the bound
(1.55) $\|P_{k}\partial_{x}\phi\|_{L^{\infty}_{t}L^{2}_{x}}\lesssim 2^{-\sigma k}c_{k}(\sigma),\ \sigma\in[0,\sigma_{1}].$
(1.56) $\sum_{k}\|P_{k}(S-Q)\|_{L^{\infty}_{t}L^{2}_{x}}^{2}\lesssim\|S_{0}-Q\|_{L^{2}_{x}}.$
Once these bounds are established, then we can extend the above local solution to a unique global solution satisfying (1.5) and (1.6) in Theorem 1.1 via standard continuity argument.
(a) Proof of (1.55) . Define the quantity
(1.57) $\Psi(T^{\prime})=\sup_{k}\sup_{\sigma\in[0,\sigma_{1}]}c_{k}(\sigma)^{-1}2^{\sigma k}\|P_{k}(\partial_{x}\phi,\bm{\psi})\|_{L^{\infty}([-T^{\prime},T^{\prime}];L^{2}_{x})}.$
We claim that
(1.58) $\text{if}\ \Psi(T^{\prime})\leq\varepsilon^{-\frac{1}{4}},\ \text{then}\ \Psi(T^{\prime})\lesssim 1.$
Under such assumptions, we have by (1.41) that
(1.59) $\|P_{k}(v,w)\|_{L^{\infty}([-T^{\prime},T^{\prime}];L^{2}_{x})}\lesssim 2^{-(\sigma+1)k}\varepsilon^{-\frac{1}{4}}c_{k}(\sigma),\ \sigma\in[0,\sigma_{1}],$
and by Proposition 1.7 that
(1.60) $\|P_{k}\bm{\psi}\|_{L^{\infty}([-T^{\prime},T^{\prime}];L^{2}_{x})}\lesssim 2^{-\sigma k}c_{k}(\sigma),\ \sigma\in[0,\sigma_{1}].$
Applying the Bony calculus and using the relation $\partial_{m}\phi=v\Re(\psi_{m})+w\Im(\psi_{m})$, we estimate
$\|P_{k}\partial_{x}\phi\|_{L^{2}_{x}}$ $\lesssim\sum_{|k_{2}-k|\leq 4}\|P_{\leq k-5}(v,w)P_{k_{2}}\bm{\psi}\|_{L^{2}_{x}}+\sum_{|k_{1}-k|\leq 4}\|P_{k_{1}}(v,w)P_{\leq k-5}\bm{\psi}\|_{L^{2}_{x}}$
$\quad+\sum^{|k_{1}-k_{2}|\leq 8}_{k_{1},k_{2}\geq k-4}\|P_{k}(P_{k_{1}}(v,w)P_{k_{2}}\bm{\psi})\|_{L^{2}_{x}}$
$\lesssim\|P_{\leq k-5}(v,w)\|_{L^{\infty}_{x}}\sum_{|k_{2}-k|\leq 4}\|P_{k_{2}}\bm{\psi}\|_{L^{2}_{x}}+\sum_{k_{2}\leq k-5}^{|k_{1}-k|\leq 4}2^{k_{2}}\|P_{k_{1}}(v,w)\|_{L^{2}_{x}}\|P_{k_{2}}\bm{\psi}\|_{L^{2}_{x}}$
$\quad+\sum^{|k_{1}-k_{2}|\leq 8}_{k_{1},k_{2}\geq k-4}2^{k}\|P_{k_{1}}(v,w)\|_{L^{2}_{x}}\|P_{k_{2}}\bm{\psi}\|_{L^{2}_{x}}$
$\lesssim 2^{-\sigma k}c_{k}(\sigma)+2^{-(\sigma+1)k}\varepsilon^{-\frac{1}{4}}c_{k}(\sigma)\sum_{k_{2}\leq k-5}2^{k_{2}}c_{k_{2}}+\sum_{k^{\prime}\geq k-10}2^{k}2^{-(\sigma+1)k^{\prime}}\varepsilon^{-\frac{1}{4}}c_{k^{\prime}}(\sigma)\cdot c_{k^{\prime}}$
$\lesssim 2^{-\sigma k}c_{k}(\sigma),$
which yields (1.58). The sums are controlled using the slow varing property of $c_{k}$ and $c_{k}(\sigma)$.
Using the local existence results and taking $\sigma_{1}=2$ in Proposition 1.7, it’s clear that there exists $T>0$ such that $\Psi(T)\lesssim 1$. A standard continuity argument based on (1.58) shows that $\Psi(T^{\prime})\lesssim 1$ for arbitarily large $T^{\prime}$, completing the proof of (1.55).
(b) Proof of (1.56). We consider the following bounds for linearized equation.
Proposition 1.8.
Given $S$ be a solution to (1.1) with small initial data $S_{0}$ satisfying $\|S_{0}\|_{\dot{H}^{1}}\ll 1$. Under the caloric gauge related to $S$, for each initial data $\psi_{\text{lin},0}\in H^{\infty}$ there exists a unique solution $\psi_{\text{lin}}\in C(\mathbb{R};H^{\infty}_{Q})$ to (1.31) satisfying
(1.61) $\sum_{k\in\mathbb{Z}}\|P_{k}\psi_{\rm{lin}}\|_{L^{2}_{x}}\lesssim\|\psi_{\rm{lin},0}\|_{L^{2}_{x}}.$
The proof of this result is identical to the proof of Proposition 1.7, since the nonlinearities in (1.31) have similar structure to those in (1.19).
We also need the following lemma from Tataru [11].
Lemma 1.9.
([11, Prop. 3.13]) Given $S_{0}^{h}\in H^{\infty}_{Q}$ with $\|S^{h}_{0}\|_{\dot{H}^{1}}\ll 1$ for $h=0,1$. There exists exists a smooth one-parameter family $S^{h}_{0}\in C^{\infty}_{h}([0,1];H^{\infty}_{Q})$ such that
(1.62) $\|S^{h}_{0}\|_{\dot{H}^{1}}\ll 1,\ \forall h\in[0,1],$
(1.63) $\int_{0}^{1}\|\partial_{h}S^{h}_{0}\|_{L^{2}}\approx\|S^{0}_{0}-S^{1}_{0}\|_{L^{2}_{x}}.$
Let $(S_{0}^{0},S_{0}^{1})=(S_{0},Q)$ in Lemma 1.9. Applying the bound (1.61) to $\psi_{\text{lin}}=\partial_{h}S^{h}$ yields
(1.64) $\sum_{k}\|P_{k}\partial_{h}S^{h}\|_{L^{\infty}_{t}L^{2}_{x}}\lesssim\|\partial_{h}S^{h}_{0}\|_{L^{2}_{x}}.$
Integrating in $h$ from 0 to $1$ gives
(1.65) $\sum_{k}\|P_{k}(S-Q)\|_{L^{\infty}_{t}L^{2}_{x}}\lesssim\|S_{0}-Q\|_{L^{2}_{x}},$
which is (1.56).
(c) Proof of continuous extension. It’s sufficient to show that $T_{Q}$ admits an unique continous extension
(1.66) $T_{Q}:H^{1}_{Q}\mapsto C(\mathbb{R};H_{Q}^{1}),$
and the extension to higher Sobolev spaces follows similarly.
Consider a sequence of solution $S^{n}$ with initial data $S_{0}^{n}\in H^{\infty}_{Q}$ and assume that $(S_{0}^{n})$ converge to $S_{0}$ in $H_{Q}^{1}$. We shall prove that $(S^{n})$ is Cauchy in $C(\mathbb{R};H_{Q}^{1})$. Repeating the arguments used to prove (1.56), we can show that
(1.67) $\limsup_{n,m\to\infty}\|S^{n}-S^{m}\|_{L^{\infty}_{t}L^{2}_{x}}\lesssim\limsup_{n,m\to\infty}\|S_{0}^{n}-S_{0}^{m}\|_{L^{2}_{x}}=0.$
Let $\{c_{k}^{n}\}$ be the frequency envelopes associated to $S_{0}^{n}$. By the convergence of $(S_{0}^{n})$ in $H_{Q}^{1}$, these envelopes also converge in $\ell^{2}$. Therefore
(1.68) $\lim_{k\to\infty}\sup_{n}\sum_{j>k}(c_{j}^{n})^{2}=0.$
Combining this with (1.55) gives
$\|S^{n}-S^{m}\|_{L^{\infty}_{t}\dot{H}^{1}_{x}}$ $\lesssim\|P_{\leq k}(S_{0}^{n}-S_{0}^{m})\|_{\dot{H}^{1}_{x}}+\|P_{>k}(S_{0}^{n},S_{0}^{m})\|_{\dot{H}^{1}_{x}}$
(1.69) $\lesssim 2^{k}\|P_{\leq k}(S_{0}^{n}-S_{0}^{m})\|_{L^{2}_{x}}+\big(\sup_{n}\sum_{j>k}(c_{j}^{n})^{2}\big)^{\frac{1}{2}}.$
Togther with (1.67) we have
(1.70) $\limsup_{n,m\to\infty}\|S^{n}-S^{m}\|_{L^{\infty}_{t}\dot{H}^{1}_{x}}\lesssim\big(\sup_{n}\sum_{j>k}(c_{j}^{n})^{2}\big)^{\frac{1}{2}}.$
Letting $k\to\infty$ and using (1.68), we obtain
(1.71) $\limsup_{n,m\to\infty}\|S^{n}-S^{m}\|_{L^{\infty}_{t}\dot{H}^{1}_{x}}=0,$
which completes the proof since $H^{1}_{Q}=\dot{H}^{1}_{x}\cap L_{Q}^{2}$. ∎
2. Preliminaries
2.1. Multilinear expression
Following Tao [10], we introduce a convenient notation for describing multi-linear expressions of product type.
Denote by $u^{y}(x):=u(x-y)$ the translation of $u(x)$. For multilinear operators, let $L$ be the integral form
(2.1) $L(u_{1},u_{2},\cdots,u_{k})(x)=\int K(y_{1},\cdots,y_{k})u_{1}^{y_{1}}(x)\cdots u_{k}^{y_{k}}(x)\mathrm{d}y,$
where $K$ can be integrable kernel or, more general, bounded measure (including product type expression, for example). The kernels may change from line to line, but we require that these kernels has uniformly bounded mass.
This $L$ notation will turn out to useful for expressing matrix coefficients, Littlewood-Paley multipliers $P_{k}$ etc., whenever these structures are not being exploited. For example, the $L$ notation is invariant under permutation of standard Littlewood-Paley operators
$L(P_{k}u_{1},u_{2},\cdots,u_{k})=L(u_{1},u_{2},\cdots,u_{k}),$
$P_{k}L(u_{1},u_{2},\cdots,u_{k})=L(u_{1},u_{2},\cdots,u_{k}).$
The same holds for $P_{>k},\ P_{<k}$, etc. Furthermore, this notation also interacts well with the composition of Littlewood-Paley operators and Riesz type operators
(2.2) $L(R_{m}P_{k}u_{1},u_{2},\cdots,u_{k})=L(P_{k}u_{1},u_{2},\cdots,u_{k}).$
Multilinear estimates are not invariant under separate translations for the factors. To obtain similar bounds for these $L$ notations, we shall allow translations in the multilinear estimates. For example, if we have the bounds with translation invariant norm $\|\cdot\|_{X}$
$\sup_{y_{\alpha}}\|u_{1}u_{2}^{y_{2}}u_{3}^{y_{3}}\|_{X}\leq C,$
then for any bilinear form $L$ with integrable kernel we have
$\|L(u_{1},u_{2},u_{3})\|_{X}$ $\leq\sup_{y_{\alpha}}\int K(y_{1},y_{2},y_{3})\|u_{1}^{y_{1}}u_{2}^{y_{2}}u_{3}^{y_{3}}\|_{X}\mathrm{d}y$
$=\sup_{y_{\alpha}}\int K(y_{1},y_{2},y_{3})\|u_{1}u_{2}^{y_{2}}u_{3}^{y_{3}}\|_{X}\mathrm{d}y\lesssim C.$
2.2. Linear estimates
Let $G=L^{4}_{t,x}\cap L^{\infty}_{t}L^{2}_{x}$ denote the standard Strichartz space. To establish Strichartz estimates for solutions to the linear ultrahyperbolic Schrödinger equation
(2.3) $(i\partial_{t}+\mu_{l}\partial_{l}^{2})u=N,\qquad\mu_{l}=\delta_{1l}-\delta_{2l},$
we follow the approach in [2] and employ adapted $U^{p}$ and $V^{p}$ spaces. These spaces provide a flexible framework for controlling the dispersive properties of the solution operator $e^{it(\partial_{1}^{2}-\partial_{2}^{2})}$.
Definition 2.1.
Let $1\leq p<\infty$, then $U^{p}_{\text{UH}}$ is an atomic space whose atoms are piecewise solutions to the linear equation, i.e.
(2.4) $u=\sum_{k}1_{[t_{k},t_{k+1})}e^{it(\partial_{1}^{2}-\partial_{2}^{2})}u_{k},\ \sum_{k}\|u_{k}\|_{L^{2}}^{p}=1.$
And we equip the $U^{p}_{\text{UH}}$ with the norm
(2.5) $\|u\|_{U^{p}_{\text{UH}}}=\inf\{\sum_{\lambda}|c_{\lambda}|\ |\ u=\sum_{\lambda}c_{\lambda}u_{\lambda},\ u_{\lambda}\ \text{are}\ U^{p}_{\text{UH}}\ \text{atoms}\}.$
The $V^{p}_{\text{UH}}$ is the space of right continuous functions $v\in L^{\infty}_{t}L_{x}^{2}$ such that
(2.6) $\|v\|^{p}_{V^{p}_{\text{UH}}}=\|v\|^{p}_{L^{\infty}_{t}L_{x}^{2}}+\sup_{\{t_{k}\}\nearrow}\sum_{k}\|e^{-it_{k}(\partial_{1}^{2}-\partial_{2}^{2})}v(t_{k})-e^{-it_{k+1}(\partial_{1}^{2}-\partial_{2}^{2})}v(t_{k+1})\|_{L^{2}},$
where the supremum is taken over increasing sequences $\{t_{k}\}$.
Theorem 2.2.
We have the following embeddings
(2.7) $U^{p}_{\text{UH}}\hookrightarrow V^{p}_{\text{UH}}\hookrightarrow U^{q}_{\text{UH}}\hookrightarrow L^{\infty}_{t}L_{x}^{2},\ 1\leq p<q<\infty.$
Let $DV^{p}_{\text{UH}}$ be the space of functions
(2.8) $DV^{p}_{\text{UH}}=\{(i\partial_{t}+\partial_{1}^{2}-\partial_{2}^{2})u\ |\ u\in V^{p}\}$
with the induced norm. And for the solution u to (2.3) we have the easy estimate:
(2.9) $\|u\|_{V^{p}_{\text{UH}}}\lesssim\|u\|_{L^{\infty}_{t}L_{x}^{2}}+\|N\|_{DV^{p}_{\text{UH}}}.$
Moreover, we have the duality relation
(2.10) $(DV^{p}_{\text{UH}})^{*}=U^{p^{\prime}}_{\text{UH}},\ \ \text{for}\ \frac{1}{p}+\frac{1}{p^{\prime}}=1.$
Finally, for $G=L^{4}_{t,x}\cap L^{\infty}_{t}L^{2}_{x}$, we have the Strichartz-type estimate
(2.11) $\|u\|_{G}\lesssim\|u\|_{L^{\infty}_{t}L^{2}_{x}}+\|N\|_{DV^{2}_{\text{UH}}}\lesssim\|u|_{t=0}\|_{L^{2}_{x}}+\|N\|_{DV^{2}_{\text{UH}}}+I(u,N)^{\frac{1}{2}},$
where $I(u,N)$ denotes the interaction term
(2.12) $I(u,N)=\sup_{y}\|u^{y}\cdot N\|_{L^{1}_{t,x}},$
which arises naturally in energy estimates.
Remark 2.3.
The spaces $U^{p}_{\text{UH}}$ and $V^{p}_{\text{UH}}$ are adaptations of the standard $U^{p}$ and $V^{p}$ spaces to the specific dispersive properties of the ultrahyperbolic Schrödinger operator $i\partial_{t}+\partial_{1}^{2}-\partial_{2}^{2}$. Their construction follows the general theory developed in [7, Chap. 4], where analogous spaces are introduced for various dispersive equations. The key properties—embeddings, duality, and connection to Strichartz estimates—are proved using the methods outlined therein.
2.3. Bilinear estimates
The following div-curl lemma, which was first introduced by the third author [14], plays a crucial role in our proof.
Lemma 2.4 (div-curl Lemma).
Suppose that $f^{ij},i,j=1,2$ satisfy
$\begin{gathered}\left\{\begin{array}[]{l}\partial_{t}f^{11}+\partial_{x}f^{12}=G^{1},\\ \partial_{t}f^{21}-\partial_{x}f^{22}=G^{2},\\ \end{array}\right.\\ f^{11},f^{12},f^{21},f^{22}\rightarrow 0,\text{ as }x\rightarrow\infty,\end{gathered}$
then it holds that
(2.13) $\int_{-\infty}^{+\infty}\int_{-\infty}^{+\infty}f^{11}f^{22}+f^{12}f^{21}\mathrm{d}x\mathrm{d}t$
$\leq$ $~2(\|f^{11}\|_{L^{\infty}_{t}L^{1}_{x}}+\|G^{1}\|_{L^{1}_{t,x}})\cdot(\|f^{21}\|_{L^{\infty}_{t}L^{1}_{x}}+\|G^{2}\|_{L^{1}_{t,x}}).$
provided that the right side is bounded.
Proof.
The same computation as in [12] yields
$\frac{\partial}{\partial t}\int_{x<y}f^{11}(t,x)f^{21}(t,y)\mathrm{d}x\mathrm{d}y+\int_{-\infty}^{+\infty}\left(f^{11}f^{21}+f^{12}f^{21}\right)\mathrm{d}x$
$=$ $\int_{-\infty}^{+\infty}\left(\int_{-\infty}^{x}f^{11}(t,y)\mathrm{d}y\right)G^{2}(t,x)\mathrm{d}x+\int_{-\infty}^{+\infty}\left(\int^{+\infty}_{x}f^{21}(t,y)\mathrm{d}y\right)G^{1}(t,x)\mathrm{d}x,$
then (2.13) follows by integrating the above inequality with respect to $t$ over $\mathbb{R}$. ∎
To proceed further we compute the conservation law of (2.3).
Multiplying (2.3) with $\bar{u}$ and taking the image part, we have the mass conservation law
(2.14) $\Im[\bar{u}(i\partial_{t}+\mu_{l}\partial_{l}^{2})u]=\frac{1}{2}\partial_{t}|u|^{2}+\mu_{l}\partial_{l}\Im(\bar{u}\partial_{l}u)=\Im(\bar{u}N).$
Multiplying (2.3) with $\partial_{m}\bar{u}$ and taking the real part, we have
$\Re[\partial_{m}\bar{u}(i\partial_{t}+\mu_{l}\partial_{l}^{2})u]$ $=-\Im(\partial_{m}\bar{u}\partial_{t}u)+\mu_{l}\Re(\partial_{m}\bar{u}\partial_{l}^{2}u)$
$=\frac{1}{2}\Im(\partial_{m}u\partial_{t}\bar{u}-\partial_{m}\bar{u}\partial_{t}u)+\mu_{l}\partial_{l}\Re(\partial_{m}\bar{u}\partial_{l}u)-\mu_{l}\Re(\partial_{m}\partial_{l}\bar{u}\partial_{l}u)$
(2.15) $=\frac{1}{2}\partial_{t}\Im(\bar{u}\partial_{m}u)-\frac{1}{2}\partial_{m}\Im(\bar{u}\partial_{t}u)+\mu_{l}\partial_{l}\Re(\partial_{m}\bar{u}\partial_{l}u)-\frac{1}{2}\mu_{l}\partial_{m}|\partial_{l}u|^{2}=\Re(\partial_{m}\bar{u}N),$
and
$\partial_{m}\Im(\bar{u}\partial_{t}u)$ $=\mu_{l}\partial_{m}\Im(i\bar{u}\partial_{l}^{2}u)-\partial_{m}\Im(i\bar{u}N)=\mu_{l}\partial_{m}\Re(\bar{u}\partial_{l}^{2}u)-\partial_{m}\Re(\bar{u}N)$
$=\mu_{l}\partial_{m}\partial_{l}\Re(\bar{u}\partial_{l}u)-\mu_{l}\partial_{m}|\partial_{l}u|^{2}-\partial_{m}\Re(\bar{u}N)$
(2.16) $=\frac{1}{2}\mu_{l}\partial_{m}\partial^{2}_{l}|u|^{2}-\mu_{l}\partial_{m}|\partial_{l}u|^{2}-\partial_{m}\Re(\bar{u}N).$
Combining above, we obtain the momentum conservation law
(2.17) $\frac{1}{2}\partial_{t}\Im(\bar{u}\partial_{m}u)+\mu_{l}\partial_{l}\Re(\partial_{m}\bar{u}\partial_{l}u)-\frac{1}{4}\mu_{l}\partial_{m}\partial^{2}_{l}|u|^{2}=\Re(\partial_{m}\bar{u}N)-\frac{1}{2}\partial_{m}\Re(\bar{u}N).$
Proposition 2.5.
Given solutions $u,v$ to the equations
(2.18) $(i\partial_{t}+\mu_{l}\partial_{l}^{2})u=N,$
(2.19) $(i\partial_{t}+\mu_{l}\partial_{l}^{2})v=N^{\prime},$
then for $k_{2}\geq k_{1}+80$ we have
(2.20) $\sup_{y}\Big\|\|P_{k_{1}}u^{y}\|_{L^{2}_{\widehat{x}_{m}}}\|\Xi_{m}(D)P_{k_{2}}v\|_{L^{2}_{\widehat{x}_{m}}}\Big\|_{L^{2}_{t,x_{m}}}$
$\lesssim 2^{-\frac{k_{2}}{2}}(\|P_{k_{1}}u\|_{L^{\infty}_{t}L^{2}_{x}}+I(P_{k_{1}}u,P_{k_{1}}N)^{\frac{1}{2}})(\|P_{k_{2}}v\|_{L^{\infty}_{t}L^{2}_{x}}+I(P_{k_{2}}v,P_{k_{2}}N^{\prime})^{\frac{1}{2}}),$
where we denote $\widehat{x}_{m}=(x_{l})_{l\neq m},\ I(u,N):=\sup_{y}\|u^{y}N\|_{L^{1}_{t,x}}$ and $\Xi_{m}(D)$ is the zero-th order Fourier multiplier supported in the Fourier region $\{|\xi|\lesssim|\xi_{m}|\}$ such that $\sum_{m=1,2}\Xi_{m}(D)=\text{Id}$.
Proof.
Applying the operator $P_{k_{1}}$ and $\Xi_{m}(D)P_{k_{2}}$ to and (2.19) respectively, we get the integrated mass conservation law of $u^{y}_{1}:=P_{k_{1}}u^{y}$ and momentum conservation law of $v_{2}:=\Xi_{m}(D)P_{k_{2}}v$ as follows:
(2.21) $\frac{1}{2}\partial_{t}\int|u^{y}_{1}|^{2}\mathrm{d}\widehat{x}_{m}+\mu_{m}\partial_{m}\int\Im(\overline{u^{y}_{1}}\partial_{m}u^{y}_{1})\mathrm{d}\widehat{x}_{m}=\int\Im(\overline{u^{y}_{1}}P_{k_{1}}N^{y})\mathrm{d}\widehat{x}_{m},$
(2.22) $\frac{1}{2}\partial_{t}\int\Im(\overline{v_{2}}\partial_{m}v_{2})\mathrm{d}\widehat{x}_{m}+\mu_{m}\partial_{m}\int\big(|\partial_{m}v_{2}|^{2}-\frac{1}{4}\partial_{m}^{2}|v_{2}|^{2}\big)\mathrm{d}\widehat{x}_{m}$
$\quad\quad=\Re(\partial_{m}\overline{v_{2}}\Xi_{m}(D)P_{k_{2}}N^{\prime})-\frac{1}{2}\partial_{m}\Re(\overline{v_{2}}\Xi_{m}(D)P_{k_{2}}N^{\prime})=2^{k_{2}}L(P_{k_{2}}v,P_{k_{2}}N^{\prime}).$
Applying the div-curl Lemma 2.4 to above conservation laws with respect to variable $(t,x_{m})$, after integration by parts we have
$\sup_{y}\iint\Big(\int|u^{y}_{1}|^{2}\mathrm{d}\widehat{x}_{m}\int|\partial_{m}v_{2}|^{2}\mathrm{d}\widehat{x}_{m}\Big)\mathrm{d}t\mathrm{d}x_{m}$
$\sup_{y}\iint\Big(\int\Re(\overline{u^{y}_{1}}\partial_{m}u^{y}_{1})\mathrm{d}\widehat{x}_{m}\int\Re(\overline{v_{2}}\partial_{m}v_{2})-\int\Im(\overline{u^{y}_{1}}\partial_{m}u^{y}_{1})\mathrm{d}\widehat{x}_{m}\int\Im(\overline{v_{2}}\partial_{m}v_{2})\mathrm{d}\widehat{x}_{m}\Big)\mathrm{d}t\mathrm{d}x_{m}$
$=:A+B\lesssim 2^{k_{2}}(\|P_{k_{1}}u\|^{2}_{L^{\infty}_{t}L^{2}_{x}}+I(P_{k_{1}}u,P_{k_{1}}N))(\|P_{k_{2}}v\|^{2}_{L^{\infty}_{t}L^{2}_{x}}+I(P_{k_{2}}v,P_{k_{2}}N^{\prime})).$
And it’s clear that
$\sup_{y}\Big\|\|P_{k_{1}}u^{y}\|_{L^{2}_{\widehat{x}_{m}}}\|\Xi_{m}(D)P_{k_{2}}v\|_{L^{2}_{\widehat{x}_{m}}}\Big\|_{L^{2}_{t,x_{m}}}^{2}$
(2.23) $\lesssim 2^{-2k_{2}}\sup_{y^{\prime}}\Big\|\|u_{1}^{y^{\prime}}\|_{L^{2}_{\widehat{x}_{m}}}\|\partial_{m}v_{2}\|_{L^{2}_{\widehat{x}_{m}}}\Big\|^{2}_{L^{2}_{t,x_{m}}}\lesssim 2^{-2k_{2}}A,$
and
$|B|$ $\lesssim\sup_{y}\iint\Big(\|u_{1}^{y}\|_{L^{2}_{\widehat{x}_{m}}}\|\partial_{m}u^{y}_{1}\|_{L^{2}_{\widehat{x}_{m}}}\|v_{2}\|_{L^{2}_{\widehat{x}_{m}}}\|\partial_{m}v_{2}\|_{L^{2}_{\widehat{x}_{m}}}\Big)\ \mathrm{d}t\mathrm{d}x_{m}$
(2.24) $\lesssim 2^{k_{1}-k_{2}}\sup_{y}\iint\Big(\|u_{1}^{y}\|^{2}_{L^{2}_{\widehat{x}_{m}}}\|\partial_{m}v_{2}\|^{2}_{L^{2}_{\widehat{x}_{m}}}\Big)\ \mathrm{d}t\mathrm{d}x_{m}\lesssim 2^{k_{1}-k_{2}}A.$
The conclusion follows from plugging the above inequalities and using the fact that $k_{2}\gg k_{1}$. ∎
Using Proposition 2.5 and Strichartz estimate, we obtain the core bilinear estimate in this paper:
Proposition 2.6.
We have
(2.25) $\sup_{y}\|P_{k_{1}}u^{y}P_{k_{2}}v\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k_{1}-k_{2}|}{2}}\mathcal{D}(P_{k_{1}}u)\mathcal{D}(P_{k_{2}}v),$
where $\mathcal{D}(u):=\|u\|_{G}+\sup_{y}\|u^{y}\cdot(i\partial_{t}+\mu_{l}\partial_{l}^{2})u\|_{L^{1}_{t,x}}$.
Proof.
If $|k_{1}-k_{2}|\leq 100$, the bound for (2.25) is straightforward; If $|k_{1}-k_{2}|\geq 100$, without loss of generality we assume $k_{2}\geq k_{1}$. We expand
$\|P_{k_{1}}u^{y}P_{k_{2}}v\|_{L^{2}_{t,x}}$ $\leq\sum_{m=1,2}\|P_{k_{1}}u\Xi_{m}(D)P_{k_{2}}v\|_{L^{2}_{t,x}}$
$\leq\sum_{m=1,2}\Big\|\|P_{k_{1}}u^{y}\|_{L^{\infty}_{\widehat{x}_{m}}}\|\Xi_{m}(D)P_{k_{2}}v\|_{L^{2}_{\widehat{x}_{m}}}\Big\|_{L^{2}_{t,x_{m}}}$
(2.26) $\lesssim 2^{\frac{k_{1}}{2}}\sum_{m=1,2}\Big\|\|P_{k_{1}}u^{y}\|_{L^{2}_{\widehat{x}_{m}}}\|\Xi_{m}(D)P_{k_{2}}v\|_{L^{2}_{\widehat{x}_{m}}}\Big\|_{L^{2}_{t,x_{m}}},$
which can be controlled by Proposition 2.5. ∎
3. Estimates in the heat flow direction
We begin with the following frequency-localized product estimate based on Bony’s paraproduct decomposition.
Lemma 3.1.
Given $f,g$ Schwartz function, let
(3.1) $\alpha_{k}(f)=\sum_{|j-k|\leq 20}\|P_{j}f\|_{G},\ \alpha_{k}(g)=\sum_{|j-k|\leq 20}\|P_{j}g\|_{G},$
then we have
(3.2) $\|P_{k}(fg)\|_{G}\lesssim\sum_{k^{\prime}\leq k}2^{k^{\prime}}[\alpha_{k^{\prime}}(f)\alpha_{k}(g)+\alpha_{k^{\prime}}(g)\alpha_{k}(f)]+\sum_{k^{\prime}\geq k}2^{k}\alpha_{k^{\prime}}(f)\alpha_{k^{\prime}}(g).$
Proof.
We apply the Bony calculus decomposition to the product fgfg
(3.3) $P_{k}(fg)=\sum_{k_{1}\leq k-4}^{|k_{2}-k|\leq 4}P_{k}(P_{k_{1}}fP_{k_{2}}g)+\sum_{k_{2}\leq k-4}^{|k_{1}-k|\leq 4}P_{k}(P_{k_{1}}fP_{k_{2}}g)+\sum_{k_{1},k_{2}\geq k-4}^{|k_{1}-k_{2}|\leq 8}P_{k}(P_{k_{1}}fP_{k_{2}}g).$
Following standard terminology, we refer to these terms respectively as the Low-High, High-Low, and High-High paraproduct interactions.
For the Low-High interaction, we use Bernstein-type inequalities to obtain
(3.4) $\|P_{k_{1}}fP_{k_{2}}g\|_{G}\lesssim\|P_{k_{1}}f\|_{L^{\infty}_{t,x}}\|P_{k_{2}}g\|_{G}\lesssim 2^{k_{1}}\alpha_{k_{1}}(f)\alpha_{k_{2}}(g).$
The High-Low case is symmetric.
For the High-High interaction, we have
(3.5) $\|P_{k_{1}}fP_{k_{2}}g\|_{G}\lesssim 2^{k}\|P_{k_{1}}fP_{k_{2}}g\|_{L^{\infty}_{t}L^{1}_{x}\cap L^{4}_{t}L^{\frac{4}{3}}_{x}}\lesssim 2^{k}\alpha_{k_{1}}(f)\alpha_{k_{2}}(g).$
The conclusion follows by summing these estimates over all frequency interactions. ∎
Lemma 3.2.
Let $f(s),g(s)$ be Schwartz functions with the norm $\alpha_{k}(f(s)),\alpha_{k}(g(s))$ defined as in Lemma 3.1 . Suppose that for $s\in[2^{2j-1},2^{2j+1}],\ n\geq 3$ and $\sigma\in[0,\sigma_{1}]$, the following bounds hold
(3.6) $\alpha_{k}(f(s))\lesssim 2^{-\sigma k}\langle 2^{2k}s\rangle^{-n}\beta_{k,j}(\sigma),\ \alpha_{k}(g(s))\lesssim 2^{-\sigma k}\langle 2^{2k}s\rangle^{-n}\eta_{k,j}(\sigma),$
(3.7) $\alpha_{k}(h(s))\lesssim 2^{-(\sigma-1)k}\langle 2^{2k}s\rangle^{-n}\rho_{k,j}(\sigma).$
where $\beta_{k,j},\ \eta_{k,j}$ are uniformly slowly varying in $k$. That is, for some sufficiently small $\widetilde{\delta}\ll 1$,
(3.8) $\beta_{k,j}(\sigma)\lesssim 2^{\widetilde{\delta}|k-k^{\prime}|}\beta_{k^{\prime},j}(\sigma),\ \forall k,k^{\prime},j\in\mathbb{Z},$
and similarly for $\eta_{k,j}(\sigma),\rho_{k,j}(\sigma)$.
Then, defining $(\beta_{k,j},\eta_{k,j},\rho_{k,j}):=(\beta_{k,j}(0),\eta_{k,j}(0),\rho_{k,j}(0))$, we have the following refined product estimates
(3.9) $\|P_{k}(f(s)g(s))\|_{G}\lesssim$ $~2^{-\sigma k}\min{\{2^{k},2^{-j}\}}^{1-2\widetilde{\delta}}2^{-2\widetilde{\delta}j}\langle 2^{2k}s\rangle^{-n}(\beta_{-j,j}\eta_{k,j}(\sigma)+\eta_{-j,j}\beta_{k,j}(\sigma)),$
(3.10) $\|P_{k}(f(s)\partial_{x}g(s))\|_{G}$ $\lesssim 2^{-\sigma k}\langle 2^{2k}s\rangle^{-n}2^{k-j}(\beta_{-j,j}(\eta_{k,j}(\sigma)+1_{\{k+j\leq 0\}}\eta_{-j,j}(\sigma))+\eta_{-j,j}\beta_{k,j}(\sigma)),$
$\lesssim 2^{-\sigma k}\langle 2^{2k}s\rangle^{-n}2^{k-j}(\langle 2^{-\frac{k+j}{4}}\rangle\beta_{-j,j}\eta_{k,j}(\sigma)+\eta_{-j,j}\beta_{k,j}(\sigma)),$
(3.11) $\|P_{k}(f(s)h(s))\|_{G}$ $\lesssim 2^{-(\sigma-1)k}\langle 2^{2k}s\rangle^{-n}2^{-j}(\rho_{-j,j}(\beta_{k,j}(\sigma)+1_{\{k+j\leq 0\}}\beta_{-j,j}(\sigma))+\beta_{-j,j}\rho_{k,j}(\sigma)).$
Moreover, for $\sigma\geq\frac{1}{5}$ the bound (3.11) can be improved to
(3.12) $\|P_{k}(f(s)h(s))\|_{G}\lesssim 2^{-(\sigma-1)k}\langle 2^{2k}s\rangle^{-n}2^{-j}(\rho_{-j,j}\beta_{k,j}(\sigma)+\beta_{-j,j}\rho_{k,j}(\sigma))$
Proof.
We begin by establishing estimates for the low-frequency sum. We compute:
$\sum_{k^{\prime}\leq k}2^{k^{\prime}}\alpha_{k^{\prime}}(f(s))\lesssim\sum_{k^{\prime}\leq k}\langle 2^{2k^{\prime}+2j}\rangle^{-n}\beta_{k^{\prime},j}$
$\leq 1_{\{k+j\leq 0\}}\sum_{k^{\prime}\leq k}2^{k^{\prime}}\beta_{k^{\prime},j}+1_{\{k+j\geq 0\}}\Big(\sum_{k^{\prime}\leq-j}2^{k^{\prime}}\beta_{k^{\prime},j}+\sum_{-j\leq k^{\prime}\leq k}2^{k^{\prime}}\langle 2^{2(k^{\prime}+j)}\rangle^{-n}\beta_{k^{\prime},j}\Big)$
$\lesssim 1_{\{k+j\leq 0\}}\beta_{-j,j}\sum_{k^{\prime}\leq k}2^{k^{\prime}}2^{-\widetilde{\delta}(k^{\prime}+j)}$
$\quad+1_{\{k+j\geq 0\}}\beta_{-j,j}\Big(\sum_{k^{\prime}\leq-j}2^{k^{\prime}}2^{-\widetilde{\delta}(j+k^{\prime})}+\sum_{-j\leq k^{\prime}\leq k}2^{k^{\prime}}\langle 2^{2(k^{\prime}+j)}\rangle^{-n}2^{\widetilde{\delta}(k^{\prime}+j)}\Big)$
(3.13) $\lesssim\min{\{2^{k},2^{-j}\}}^{1-\widetilde{\delta}}2^{-\widetilde{\delta}j}\beta_{-j,j},$
where we use the slow variation property of $\beta_{k,j}$. Similarly we have
(3.14) $\sum_{k^{\prime}\leq k}2^{k^{\prime}}\alpha_{k^{\prime}}(f(s))\lesssim\min{\{2^{k},2^{-j}\}}^{1-\widetilde{\delta}}2^{-\widetilde{\delta}j}\eta_{-j,j},$
(3.15) $\sum_{k^{\prime}\leq k}2^{k^{\prime}}\alpha_{k^{\prime}}(h(s))\lesssim 2^{k}\min{\{2^{k},2^{-j}\}}^{1-\widetilde{\delta}}2^{-\widetilde{\delta}j}\rho_{-j,j}$
We now apply Lemma 3.1 to estimate the product terms. The estimates split naturally into High-Low, Low-High, and High-High frequency interactions according to the paraproduct decomposition.
High-Low and Low-High Interactions: Consider the sums over $k_{1},k_{2}$ where one frequency is much lower than $k$. For the product $P_{k}(P_{k_{1}}f(s)P_{k_{2}}g(s))$, we have
$\sum_{k_{1}\leq k-4}^{|k_{2}-k|\leq 4}+\sum_{k_{2}\leq k-4}^{|k_{1}-k|\leq 4}\|P_{k}(P_{k_{1}}f(s)P_{k_{2}}g(s))\|_{G}$
$\lesssim\alpha_{k}(g(s))\sum_{k^{\prime}\leq k}2^{k^{\prime}}\alpha_{k^{\prime}}(f(s))+\alpha_{k}(f(s))\sum_{k^{\prime}\leq k}2^{k^{\prime}}\alpha_{k^{\prime}}(g(s))$
(3.16) $\lesssim 2^{-\sigma k}\min{\{2^{k},2^{-j}\}}^{1-\widetilde{\delta}}2^{-\widetilde{\delta}j}\langle 2^{2k}s\rangle^{-n}(\beta_{-j,j}\eta_{k,j}(\sigma)+\eta_{-j,j}\beta_{k,j}(\sigma)).$
For the product $P_{k}(P_{k_{1}}f(s)P_{k_{2}}\partial_{x}g(s))$, the additional derivative contributes a factor of $2^{k}$, which leads to
$\sum_{k_{1}\leq k-4}^{|k_{2}-k|\leq 4}+\sum_{k_{2}\leq k-4}^{|k_{1}-k|\leq 4}\|P_{k}(P_{k_{1}}f(s)P_{k_{2}}\partial_{x}g(s))\|_{G}$
$\lesssim 2^{k}\alpha_{k}(g(s))\sum_{k^{\prime}\leq k}2^{k^{\prime}}\alpha_{k^{\prime}}(f(s))+2^{k}\alpha_{k}(f(s))\sum_{k^{\prime}\leq k}2^{k^{\prime}}\alpha_{k^{\prime}}(g(s))$
(3.17) $\lesssim 2^{-\sigma k}2^{k-j}\langle 2^{2k}s\rangle^{-n}(\beta_{-j,j}\eta_{k,j}(\sigma)+\eta_{-j,j}\beta_{k,j}(\sigma)).$
For the product $P_{k}(P_{k_{1}}f(s)P_{k_{2}}h(s))$, similar computation shows that
$\sum_{k_{1}\leq k-4}^{|k_{2}-k|\leq 4}+\sum_{k_{2}\leq k-4}^{|k_{1}-k|\leq 4}\|P_{k}(P_{k_{1}}f(s)P_{k_{2}}h(s))\|_{G}$
$\lesssim\alpha_{k}(h(s))\sum_{k^{\prime}\leq k}2^{k^{\prime}}\alpha_{k^{\prime}}(f(s))+\alpha_{k}(f(s))\sum_{k^{\prime}\leq k}2^{k^{\prime}}\alpha_{k^{\prime}}(h(s))$
(3.18) $\lesssim 2^{-(\sigma-1)k}\min{\{2^{k},2^{-j}\}}^{1-\widetilde{\delta}}2^{-\widetilde{\delta}j}\langle 2^{2k}s\rangle^{-n}(\beta_{-j,j}\rho_{k,j}(\sigma)+\rho_{-j,j}\beta_{k,j}(\sigma)).$
which is consistent with the bound (3.11) and (3.12).
High-High Interactions: It remains to estimate the sum where both frequencies are high and comparable. For the product $P_{k}(P_{k_{1}}f(s)P_{k_{2}}g(s))$, we have
$\sum_{k_{1},k_{2}\geq k-4}^{|k_{1}-k_{2}|\leq 8}\|P_{k}(P_{k_{1}}f(s)P_{k_{2}}g(s))\|_{G}\lesssim\sum_{k^{\prime}\geq k}2^{k}\alpha_{k^{\prime}}(f(s))\alpha_{k^{\prime}}(g(s))$
$\lesssim\sum_{k^{\prime}\geq k}2^{k}2^{-\sigma k^{\prime}}\langle 2^{2k^{\prime}+2j}\rangle^{-2n}\beta_{k^{\prime},j}\eta_{k^{\prime},j}(\sigma)$
$\lesssim 1_{\{k+j\geq 0\}}\sum_{k^{\prime}\geq k}\ldots+1_{\{k+j\leq 0\}}\Big(\sum_{k\leq k^{\prime}\leq-j}\ldots+\sum_{k^{\prime}\geq-j}\ldots\Big)$
$\lesssim 1_{\{k+j\geq 0\}}2^{-j}2^{-\sigma k}\beta_{-j,j}\eta_{k,j}(\sigma)\sum_{k^{\prime}\geq k}2^{k+j}2^{\widetilde{\delta}(k^{\prime}-k)}2^{\widetilde{\delta}(k^{\prime}+j)}\langle 2^{2k^{\prime}+2j}\rangle^{-2n}$
$\quad+1_{\{k+j\leq 0\}}2^{-\sigma k}\beta_{-j,j}\eta_{k,j}(\sigma)$
$\qquad\cdot 2^{k}\Big(\sum_{k\leq k^{\prime}\leq-j}2^{2\widetilde{\delta}(k^{\prime}-k)}2^{-\widetilde{\delta}(k^{\prime}+j)}+\sum_{k^{\prime}\geq-j}2^{\widetilde{\delta}[(k^{\prime}-k)+(k^{\prime}+j)]}\langle 2^{2k^{\prime}+2j}\rangle^{-2n}\Big)$
(3.19) $\lesssim 1_{\{k+j\geq 0\}}2^{-\sigma k}2^{-j}\langle 2^{2k+2j}\rangle^{2-2n}\beta_{-j,j}\eta_{k,j}(\sigma)+1_{\{k+j\leq 0\}}2^{-\sigma k}2^{(1-2\widetilde{\delta})k-2\widetilde{\delta}j}\beta_{-j,j}\eta_{k,j}(\sigma),$
which is consistent with the stated bound (3.9).
For $P_{k}(P_{k_{1}}f(s)P_{k_{2}}\partial_{x}g(s))$ in the High-High regime, the derivative gives an extra factor of $2^{k^{\prime}}$, which leads to
$\sum_{k_{1},k_{2}\geq k-4}^{|k_{1}-k_{2}|\leq 8}\|P_{k}(P_{k_{1}}f(s)P_{k_{2}}\partial_{x}g(s))\|_{G}\lesssim\sum_{k^{\prime}\geq k}2^{k+k^{\prime}}\alpha_{k^{\prime}}(f(s))\alpha_{k^{\prime}}(g(s))$
$\lesssim\sum_{k^{\prime}\geq k}2^{k+k^{\prime}}2^{-\sigma k^{\prime}}\langle 2^{2k^{\prime}+2j}\rangle^{-2n}\beta_{k^{\prime},j}\eta_{k^{\prime},j}(\sigma)$
$\lesssim 1_{\{k+j\geq 0\}}\sum_{k^{\prime}\geq k}\ldots+1_{\{k+j\leq 0\}}\Big(\sum_{k\leq k^{\prime}\leq-j}\ldots+\sum_{k^{\prime}\geq-j}\ldots\Big)$
$\lesssim 1_{\{k+j\geq 0\}}2^{k-j}2^{-\sigma k}\beta_{-j,j}\eta_{k,j}(\sigma)\sum_{k^{\prime}\geq k}2^{k^{\prime}+j}2^{\widetilde{\delta}(k^{\prime}-k)}2^{\widetilde{\delta}(k^{\prime}+j)}\langle 2^{2k^{\prime}+2j}\rangle^{-2n}$
$\quad+1_{\{k+j\leq 0\}}2^{-\sigma k}\beta_{-j,j}\eta_{-j,j}(\sigma)$
$\qquad\cdot 2^{k-j}\Big(\sum_{k\leq k^{\prime}\leq-j}2^{k^{\prime}+j}2^{-2\widetilde{\delta}(k^{\prime}+j)}+\sum_{k^{\prime}\geq-j}2^{k^{\prime}+j}2^{2\widetilde{\delta}(k^{\prime}+j)}\langle 2^{2k^{\prime}+2j}\rangle^{-2n}\Big)$
$\lesssim 1_{\{k+j\geq 0\}}2^{k-j}2^{-\sigma k}\langle 2^{2k+2j}\rangle^{2-2n}\beta_{-j,j}\eta_{k,j}(\sigma)+1_{\{k+j\leq 0\}}2^{k-j}2^{-\sigma k}\beta_{-j,j}\eta_{-j,j}(\sigma),$
(3.20) $\lesssim 1_{\{k+j\geq 0\}}\ldots+1_{\{k+j\leq 0\}}2^{k-j}2^{-\sigma k}2^{-\widetilde{\delta}(k+j)}\beta_{-j,j}\eta_{k,j}(\sigma),$
matching the stated bound (3.10).
For the $P_{k}(P_{k_{1}}f(s)P_{k_{2}}h(s))$ in the High-High regime, similar computation as in (3.20) shows that
$\sum_{k_{1},k_{2}\geq k-4}^{|k_{1}-k_{2}|\leq 8}\|P_{k}(P_{k_{1}}f(s)P_{k_{2}}h(s))\|_{G}\lesssim\sum_{k^{\prime}\geq k}2^{k}\alpha_{k^{\prime}}(f(s))\alpha_{k^{\prime}}(h(s))$
$\lesssim\sum_{k^{\prime}\geq k}2^{k+k^{\prime}}2^{-\sigma k^{\prime}}\langle 2^{2k^{\prime}+2j}\rangle^{-2n}\beta_{k^{\prime},j}(\sigma)\rho_{k^{\prime},j}$
$\lesssim 1_{\{k+j\geq 0\}}2^{-j}2^{-(\sigma-1)k}\langle 2^{2k+2j}\rangle^{2-2n}\rho_{-j,j}\beta_{k,j}(\sigma)$
(3.21) $\quad\quad+1_{\{k+j\leq 0\}}2^{-j}2^{-(\sigma-1)k}\rho_{-j,j}\beta_{-j,j}(\sigma).$
To obtain (3.12) for $\sigma\geq\frac{1}{5}$, in view of (3.18) and (3.21) it’s sufficient to improve the bound in the High-High interaction when $k+j\leq 0$. We have
$1_{\{k+j\leq 0\}}\sum_{k_{1},k_{2}\geq k-4}^{|k_{1}-k_{2}|\leq 8}\|P_{k}(P_{k_{1}}f(s)P_{k_{2}}\partial_{x}h(s))\|_{G}\lesssim 1_{\{k+j\leq 0\}}\sum_{k^{\prime}\geq k}2^{k}\alpha_{k^{\prime}}(f(s))\alpha_{k^{\prime}}(h(s))$
$\lesssim 1_{\{k+j\leq 0\}}\sum_{k^{\prime}\geq k}2^{k+k^{\prime}}2^{-\sigma k^{\prime}}\langle 2^{2k^{\prime}+2j}\rangle^{-2n}\beta_{k^{\prime},j}(\sigma)\rho_{k^{\prime},j}$
$\lesssim 1_{\{k+j\leq 0\}}\beta_{-j,j}(\sigma)\rho_{-j,j}2^{k}\Big(\sum_{k^{\prime}\geq-j}2^{(1-\sigma)k^{\prime}}\langle 2^{2k^{\prime}+2j}\rangle^{-2n}2^{2\widetilde{\delta}(k^{\prime}+j)}+\sum_{k\leq k^{\prime}\leq-j}2^{(1-\sigma)k^{\prime}}2^{-2\widetilde{\delta}(k^{\prime}+j)}\Big)$
$\lesssim 1_{\{k+j\leq 0\}}\beta_{-j,j}(\sigma)\rho_{-j,j}2^{k}(\max\{2^{-(1-\sigma)j},2^{(1-\sigma)k}2^{-2\widetilde{\delta}(k+j)}\})$
$\lesssim 1_{\{k+j\leq 0\}}\beta_{-j,j}(\sigma)\rho_{-j,j}2^{k}2^{-(1-\sigma)j}\langle 2^{(1-\sigma-2\widetilde{\delta})(k+j)}\rangle$
$\lesssim 1_{\{k+j\leq 0\}}2^{\sigma(k+j)}\langle 2^{(1-\sigma-2\widetilde{\delta})(k+j)}\rangle 2^{-j}2^{-(\sigma-1)k}\rho_{-j,j}\beta_{-j,j}(\sigma)$
$\lesssim 1_{\{k+j\leq 0\}}2^{\frac{1}{5}(k+j)}2^{-j}2^{-(\sigma-1)k}\rho_{-j,j}\beta_{-j,j}(\sigma)\lesssim 1_{\{k+j\leq 0\}}2^{-j}2^{-(\sigma-1)k}\beta_{k,j}(\sigma)\rho_{-j,j},$
which sufficients to prove the bound (3.12).
This completes the proof. ∎
We now apply these technical lemmas to obtain estimates for the differentiated fields and connection coefficients $\psi_{\alpha},A_{\alpha}$ in the heat flow direction.
By (1.45) we have the following bound for $\psi_{\alpha}$ when $s\in[2^{2j-1},2^{2j+1}]$:
(3.22) $\|P_{k}\bm{\psi}(s)\|_{G}\leq 2^{-\sigma k}\langle 2^{2k}s\rangle^{-4}a_{k}(\sigma),\ \text{for}\ \sigma\in[0,\sigma_{1}],$
(3.23) $\|P_{k}\psi_{t}(s)\|_{G}\leq 2^{-(\sigma-1)k}\langle 2^{2k}s\rangle^{-4}a_{k}(\sigma),\ \text{for}\ \sigma\in[\frac{1}{5},\sigma_{1}].$
Define $B_{1}$ as the smallest constant in $1,\infty)$ such that
(3.24) $\|\bm{A}(s)\|_{G}\lesssim B_{1}2^{-\sigma k}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{k,j}(\sigma),$
where $a_{k,j}(\sigma)$ is defined by
(3.25) $a_{k,j}(\sigma):=a_{\min(-j,k)}a_{k}(\sigma).$
We note that $a_{k,j}(\sigma)$ is uniformly slowly varying in $k$. Using the slow variation of $a_{k}$, we have the lower bound
(3.26) $\langle 2^{-\delta(k+j)}\rangle^{-1}a_{-j}a_{k}(\sigma)\lesssim 1_{\{k+j\leq 0\}}a_{k}a_{k}(\sigma)+1_{\{k+j\geq 0\}}a_{-j}a_{k}(\sigma)=a_{k,j}(\sigma)$
and the upper bound
(3.27) $a_{k,j}(\sigma)\lesssim\langle 2^{-\delta(k+j)}\rangle a_{-j}a_{k}(\sigma).$
And the uniform slow variation of $a_{k,j}(\sigma)$ follows from that of $a_{k}(\sigma)$.
With these preliminaries, we apply Lemma 3.2 to control nonlinear interactions involving $\psi_{\alpha}$ and $A_{\alpha}$.
Lemma 3.3.
Let $s\in[2^{2j-1},2^{2j+1}],\sigma\in[0,\sigma_{1}]$ and assume the bound (3.24) holds. Then we have the following bilinear and trilinear estimates
(3.28) $\|P_{k}(\bm{\psi}(s)\cdot\bm{\psi}(s))\|_{G}\lesssim 2^{-\sigma k}\min{\{2^{k},2^{-j}\}}^{1-2\delta}2^{-2\delta j}\langle 2^{2k}s\rangle^{-4}a_{-j}a_{k}(\sigma),$
(3.29) $\|P_{k}(\bm{\psi}(s)\cdot\partial_{x}\bm{\psi}(s))\|_{G}\lesssim 2^{-\sigma k}\langle 2^{2k}s\rangle^{-4}2^{k-j}\langle(2^{2k}s)^{-\frac{1}{8}}\rangle a_{-j}a_{k}(\sigma),$
(3.30) $\|P_{k}(\bm{A}(s)\cdot\bm{\psi}(s))\|_{G}\lesssim B_{1}2^{-\sigma k}\min{\{2^{k},2^{-j}\}}^{1-6\delta}2^{-6\delta j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}^{2}a_{k}(\sigma),$
(3.31) $\|P_{k}(\bm{A}(s)\cdot\bm{A}(s))\|_{G}\lesssim B^{2}_{1}2^{-\sigma k}\min{\{2^{k},2^{-j}\}}^{1-6\delta}2^{-6\delta j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}^{3}a_{k}(\sigma),$
(3.32) $\|P_{k}(\bm{A}(s)\cdot\partial_{x}\bm{\psi}(s))\|_{G}\lesssim B_{1}2^{-\sigma k}2^{k-j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}\langle(2^{2k}s)^{-\frac{1}{8}}\rangle a_{-j}^{2}a_{k}(\sigma),$
(3.33) $\|P_{k}(\bm{A}(s)\cdot\psi_{t}(s))\|_{G}\lesssim B_{1}2^{-\sigma k}2^{k-j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}\langle(2^{2k}s)^{-\frac{1}{8}}\rangle a_{-j}^{2}a_{k}(\sigma),\ \sigma\geq\frac{1}{5},$
(3.34) $\|P_{k}(\bm{A}(s)\cdot\bm{A}(s)\cdot\bm{\psi}(s))\|_{G}+\|P_{k}(\bm{\psi}(s)\cdot\bm{\psi}(s)\cdot\bm{\psi}(s))\|_{G}$
$\qquad\quad\lesssim B^{2}_{1}2^{-\sigma k}2^{-2j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}(a_{-j}^{4}+a_{-j}^{2})a_{k}(\sigma),$
(3.35) $\|P_{k}(\bm{A}(s)\cdot\bm{\psi}(s)\cdot\bm{\psi}(s))\|_{G}$
$\qquad\quad\lesssim B_{1}2^{-\sigma k}(2^{2k}2^{-\frac{3(k+j)}{2}}1_{\{k+j\leq 0\}}+2^{-2j}1_{\{k+j\geq 0\}})\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}^{3}a_{k}(\sigma).$
Proof.
Bounds (3.28) and (3.29) follow directly from applying Lemma 3.2 to the product involving bound (3.22).
To obtain bounds (3.30)–(3.32), we apply Lemma 3.2 to the product forms of the bounds (3.22), (3.23) and (3.24), obtaining
(3.36) $\|P_{k}(\bm{A}(s)\cdot\bm{\psi}(s))\|_{G}$
$\quad\lesssim B_{1}2^{-\sigma k}\min{\{2^{k},2^{-j}\}}^{1-4\delta}2^{-4\delta j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}(a_{k,j}(\sigma)+a_{-j}a_{k}(\sigma))$
(3.37) $\|P_{k}(\bm{A}(s)\cdot\bm{A}(s))\|_{G}\lesssim B^{2}_{1}2^{-\sigma k}\min{\{2^{k},2^{-j}\}}^{1-4\delta}2^{-4\delta j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}^{2}a_{k,j}(\sigma),$
(3.38) $\|P_{k}(\bm{A}(s)\cdot\partial_{x}\bm{\psi}(s))\|_{G}+\|P_{k}(\bm{A}(s)\cdot\psi_{t}(s))\|_{G}$
$\quad\lesssim B_{1}2^{-\sigma k}2^{k-j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}(\langle 2^{-\frac{k+j}{4}}\rangle a_{-j}a_{k}(\sigma)+a_{k,j}(\sigma)).$
Then (3.30) and (3.31) follow from (3.36), (3.37) together with the inequality
(3.39) $\min\{2^{k},2^{-j}\}^{2\delta}a_{k,j}(\sigma)\lesssim\min\{2^{k},2^{-j}\}^{2\delta}\langle 2^{-2\delta(k+j)}\rangle a_{-j}a_{k}(\sigma)\lesssim 2^{-2\delta j}a_{-j}a_{k}(\sigma),$
while (3.32) follows from (3.38) and the upper bound (3.27) for $a_{k,j}(\sigma)$.
The bound (3.34) is obtained by applying Lemma 3.2 to (3.22), (3.28), and (3.31) while replacing the factor $\min\{2^{k},2^{-j}\}$ with $2^{-j}$ .
For (3.35), we apply Lemma 3.2 to the product $\bm{A}(s)\cdot(\bm{\psi}(s)\cdot\bm{\psi}(s))$. Using (3.28) and the lower bound (3.26) for $a_{k,j}(\sigma)$, we have
$\|P_{k}(\bm{\psi}(s)\cdot\bm{\psi}(s))\|_{G}$ $\lesssim 2^{-\sigma k}\min{\{2^{k},2^{-j}\}}^{1-2\delta}2^{-2\delta j}\langle 2^{2k}s\rangle^{-4}a_{-j}a_{k}(\sigma)$
$\lesssim 2^{-\sigma k}\min{\{2^{k},2^{-j}\}}^{1-2\delta}2^{-2\delta j}\langle 2^{2k}s\rangle^{-4}\langle 2^{-2\delta(k+j)}\rangle a_{k,j}(\sigma)$
(3.40) $\lesssim 2^{-\sigma k}2^{-j}\langle 2^{2k}s\rangle^{-4}a_{k,j}(\sigma).$
Combining this with (3.24) and (3.9) yields
$\|P_{k}(\bm{A}(s)\cdot\bm{\psi}(s)\cdot\bm{\psi}(s))\|_{G}$
$\lesssim B_{1}2^{-\sigma k}\min{\{2^{k},2^{-j}\}}^{1-4\delta}2^{-4\delta j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}\cdot 2^{-j}a_{-j,j}a_{k,j}(\sigma)$
$\lesssim B_{1}2^{-\sigma k}2^{-j}\min{\{2^{k},2^{-j}\}}^{1-4\delta}2^{-4\delta j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}^{2}a_{k,j}(\sigma)$
$\lesssim B_{1}2^{-\sigma k}2^{-j}\min{\{2^{k},2^{-j}\}}^{1-6\delta}2^{-6\delta j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}^{3}a_{k}(\sigma)$
(3.41) $\lesssim B_{1}2^{-\sigma k}(2^{2k}2^{-\frac{3(k+j)}{2}}1_{\{k+j\leq 0\}}+2^{-2j}1_{\{k+j\geq 0\}})\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}^{3}a_{k}(\sigma).$
This completes the proof. ∎
We now establish the main estimate for the connection coefficient $\bm{A}(s)$.
Proposition 3.4.
Let $s\in[2^{2j_{0}-1},2^{2j_{0}+1}],\sigma\in[0,\sigma_{1}]$. Under the bootstrap assumptions in Proposition 1.6 we have
(3.42) $\|\bm{A}(s)\|_{G}\lesssim 2^{-\sigma k}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{k,j_{0}}(\sigma).$
Proof.
We start from the identity (1.27)
(3.43) $A_{m}(s)=-\int_{s}^{+\infty}\Im(\overline{\psi_{m}}D_{l}\psi_{l})(r)\mathrm{d}r,\ m=1,2.$
Applying $P_{k}$ to (3.43) yields
(3.44) $\|P_{k}\bm{A}(s)\|_{G}\lesssim$ $~\int_{s}^{+\infty}\|P_{k}(\bm{\psi}(r)\cdot\partial_{x}\bm{\psi}(r))\|_{G}+\|P_{k}(\bm{A}(r)\cdot\bm{\psi}(r)\cdot\bm{\psi}(r))\|_{G}\mathrm{d}r.$
For the first term, using (3.29) we compute
$\int_{s}^{+\infty}\|P_{k}(\bm{\psi}(r)\cdot\partial_{x}\bm{\psi}(r))\|_{G}\mathrm{d}r$
$\lesssim\sum_{j\geq j_{0}}2^{2j}\cdot 2^{-\sigma k}\langle 2^{2k+2j}\rangle^{-4}2^{k-j}\langle 2^{-\frac{k+j}{4}}\rangle a_{-j}a_{k}(\sigma)$
$\lesssim 1_{\{k+j_{0}\geq 0\}}\sum_{j\geq j_{0}}\ldots+1_{\{k+j_{0}\leq 0\}}\Big(\sum_{j_{0}\leq j\leq-k}\ldots+\sum_{j\geq-k}\ldots\Big)$
$\lesssim 1_{\{k+j_{0}\geq 0\}}2^{-\sigma k}a_{-j_{0}}a_{k}(\sigma)\sum_{j\geq j_{0}}\langle 2^{2k+2j}\rangle^{-4}2^{k+j}2^{\delta(j-j_{0})}$
$\ +1_{\{k+j_{0}\leq 0\}}2^{-\sigma k}a_{k}a_{k}(\sigma)\Big(\sum_{j_{0}\leq j\leq-k}2^{k+j}2^{-(\frac{1}{4}+\delta)(k+j)}+\sum_{j\geq-k}\langle 2^{2k+2j}\rangle^{-4}2^{k+j}2^{\delta(k+j)}\Big)$
$\lesssim 1_{\{k+j_{0}\geq 0\}}2^{-\sigma k}\langle 2^{2k+2j}\rangle^{-\frac{7}{2}}a_{-j_{0}}a_{k}(\sigma)+1_{\{k+j_{0}\leq 0\}}2^{-\sigma k}a_{k}a_{k}(\sigma)$
(3.45) $\lesssim 2^{-\sigma k}\langle 2^{2k+2j}\rangle^{-\frac{7}{2}}a_{k,j_{0}}(\sigma).$
For the second term, using (3.35) and the bootstrap assumption $\sum_{k}a_{k}^{2}\leq\varepsilon$,
$\int_{s}^{+\infty}\|P_{k}(\bm{A}(r)\cdot\bm{\psi}(r)\cdot\bm{\psi}(r))\|_{G}\mathrm{d}r$
$\lesssim\sum_{j\geq j_{0}}B_{1}2^{-\sigma k}(2^{\frac{k+j}{2}}1_{\{k+j\leq 0\}}+1_{\{k+j\geq 0\}})\langle 2^{2k+2j}\rangle^{-\frac{7}{2}}a_{-j}^{3}a_{k}(\sigma)$
$\lesssim 1_{\{k+j_{0}\leq 0\}}B_{1}\varepsilon 2^{-\sigma k}a_{k}a_{k}(\sigma)\big(\sum_{j\geq-k}\langle 2^{2k+2j}\rangle^{-\frac{7}{2}}2^{\delta(k+j)}+\sum_{j_{0}\leq j\leq-k}2^{k+j}2^{-(\frac{1}{4}+2\delta)(k+j)}\big)$
$\quad+1_{\{k+j_{0}\geq 0\}}B_{1}\varepsilon 2^{-\sigma k}\sum_{j\geq j_{0}}\langle 2^{2k+2j}\rangle^{-\frac{7}{2}}2^{\delta(j-j_{0})}a_{-j_{0}}a_{k}(\sigma)$
$\lesssim B_{1}\varepsilon 2^{-\sigma k}(1_{\{k+j_{0}\leq 0\}}a_{k}a_{k}(\sigma)+1_{\{k+j_{0}\leq 0\}}\langle 2^{2k+2j_{0}}\rangle^{-\frac{7}{2}}a_{-j_{0}}a_{k}(\sigma))$
(3.46) $\lesssim B_{1}\varepsilon 2^{-\sigma k}\langle 2^{2k+2j_{0}}\rangle^{-\frac{7}{2}}a_{k,j_{0}}(\sigma),$
Combining (3.45) and (3.46), we have
(3.47) $\|A(s)\|_{G}\lesssim\langle B_{1}\varepsilon\rangle 2^{-\sigma k}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{k,j_{0}}(\sigma),$
which implies $B_{1}\lesssim 1+B_{1}\varepsilon$, hence $B_{1}\lesssim 1$ since $\varepsilon$ is sufficiently small. ∎
We next establish the bound for $\bm{\psi}(s)$ and $\psi_{t}(s)$ using the heat equation.
Proposition 3.5.
Under the bootstrap assumptions in Proposition 1.6, we have
(3.48) $\|P_{k}\bm{\psi}(s)\|_{G}\lesssim 2^{-\sigma k}\langle 2^{2k}s\rangle^{-4}b_{k}(\sigma),\ \text{for}\ \sigma\in[0,\sigma_{1}],$
(3.49) $\|P_{k}\psi_{t}(s)\|_{G}\lesssim 2^{k}\langle 2^{2k}s\rangle^{-4}\varepsilon^{\frac{1}{2}},\ 2^{-(\sigma-1)k}\langle 2^{2k}s\rangle^{-4}b_{k}(\sigma),\ \text{for}\ \sigma\in[\frac{1}{5},\sigma_{1}].$
Proof.
We use the heat equation for $\psi_{\alpha}$:
(3.50) $(\partial_{s}-\Delta_{x})\psi_{\alpha}=K(\psi_{\alpha}),$
where $~K(\psi)=2iA_{l}\partial_{l}\psi-(A_{l}^{2}-i\partial_{l}A_{l})\psi+i\Im(\psi\overline{\psi_{l}})\psi_{l},$
which gives the Duhamel formula
(3.51) $\psi_{\alpha}(s)=e^{s\Delta_{x}}\psi_{\alpha}(0)+\int_{0}^{s}e^{(s-r)\Delta_{x}}K(\psi_{\alpha})\mathrm{d}r.$
For $\bm{\psi}$, we rewrite $K(\bm{\psi})$ as
(3.52) $K(\bm{\psi})=\partial_{x}(\bm{A}\bm{\psi})+\bm{A}\partial_{x}\bm{\psi}+(\bm{A}\cdot\bm{A}+\bm{\psi}\cdot\bm{\psi})\psi.$
Using (3.30), (3.32), (3.34) and the bootstrap assumption $\sum_{k}a_{k}^{2}\leq\varepsilon$, for $r\in[2^{2j-1},2^{2j+1}]$ we obtain
(3.53) $\|P_{k}K(\bm{\psi})(s)\|_{G}\lesssim 2^{-\sigma k}(2^{-2j}a_{-j}^{2}+\varepsilon 2^{k-j}\langle 2^{-\frac{k+j}{2}}\rangle)\langle 2^{2k+2j}\rangle^{-\frac{7}{2}}a_{k}(\sigma).$
Assume $s\in[2^{2k_{0}-1},2^{2k_{0}+1}]$. For $k+k_{0}\geq 0$, splitting the time integral at $s/2$, we have
$\Big\|\int_{0}^{s}e^{(s-r)\Delta_{x}}P_{k}K(\psi_{\alpha})\mathrm{d}r\Big\|_{G}$
$\lesssim\int_{0}^{\frac{s}{2}}\langle 2^{2k}s\rangle^{-N}\|P_{k}K(\psi_{\alpha})(r)\|_{G}\mathrm{d}r+\int_{\frac{s}{2}}^{s}\langle 2^{2k}(s-r)\rangle^{-N}\|P_{k}K(\psi_{\alpha})(r)\|_{G}\mathrm{d}r$
$\lesssim 2^{-\sigma k}\langle 2^{2k}s\rangle^{-N}a_{k}(\sigma)\sum_{j\leq k_{0}}(a_{-j}^{2}+\varepsilon 2^{k+j}\langle 2^{-\frac{k+j}{2}}\rangle)\langle 2^{2k+2j}\rangle^{-\frac{7}{2}}$
$\quad+\varepsilon 2^{-\sigma k}(2^{-2k-2k_{0}}+2^{-k-k_{0}})\langle 2^{2k+2k_{0}}\rangle^{-\frac{7}{2}}a_{k}(\sigma)$
(3.54) $\lesssim\varepsilon 2^{-\sigma k}\langle 2^{2k+2k_{0}}\rangle^{-4}a_{k}(\sigma).$
For $k+k_{0}\leq 0$, we integrate directly
$\Big\|\int_{0}^{s}e^{(s-r)\Delta_{x}}P_{k}K(\psi_{\alpha})\mathrm{d}r\Big\|_{G}\lesssim\sum_{j\leq k_{0}}\int_{2^{2j-1}}^{2^{2j+1}}\|P_{k}K(\psi_{\alpha})(r)\|_{G}\mathrm{d}r$
(3.55) $\lesssim 2^{-\sigma k}a_{k}(\sigma)\sum_{j\leq k_{0}}(a_{-j}^{2}+\varepsilon 2^{\frac{k+j}{2}})\lesssim\varepsilon 2^{-\sigma k}a_{k}(\sigma).$
Combining both cases yields
(3.56) $\Big\|\int_{0}^{s}e^{(s-r)\Delta_{x}}P_{k}K(\psi_{\alpha})\mathrm{d}r\Big\|_{G}\lesssim\varepsilon 2^{-\sigma k}\langle 2^{2k}s\rangle^{-4}a_{k}(\sigma).$
From (3.51) and the definition of $b_{k}(\sigma)$,
$\|\bm{\psi}(s)\|_{G}$ $\lesssim\|e^{s\Delta_{x}}\bm{\psi}(0)\|_{G}+\Big\|\int_{0}^{s}e^{(s-r)\Delta_{x}}P_{k}K(\bm{\psi})\mathrm{d}r\Big\|_{G}$
$\lesssim 2^{-\sigma k}(\langle 2^{2k}s\rangle^{-N}b_{k}(\sigma)+\varepsilon\langle 2^{2k}s\rangle^{-4}a_{k}(\sigma))$
(3.57) $\lesssim 2^{-\sigma k}\langle 2^{2k}s\rangle^{-4}(b_{k}(\sigma)+\varepsilon a_{k}(\sigma)).$
For $\psi_{t}(s)$, let $B_{2}$ be the smallest number in $1,\infty)$ such that
(3.58) $\|\psi_{t}(s)\|_{G}\lesssim B_{2}\varepsilon^{\frac{1}{2}}2^{k}\langle 2^{2k}s\rangle^{-4},$
we rewrite $K(\psi_{t})$ as
(3.59) $K(\psi_{t})=\partial_{x}(\bm{A}\psi_{t})+[\partial_{x}\bm{A}+(\bm{A}\cdot\bm{A}+\bm{\psi}\cdot\bm{\psi})]\psi_{t}.$
Using (3.24), (3.27), (3.28), (3.31) and the slow varying of $a_{k}(\sigma)$, we have
(3.60) $\|P_{k}(\bm{A}(s)\cdot\bm{A}(s),\ \bm{\psi}(s)\cdot\bm{\psi}(s),\ \partial_{x}\bm{A}(s))\|_{G}$ $\lesssim 2^{-j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}^{2}.$
(3.61) or $\lesssim 2^{-\sigma k}2^{-j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}a_{k}(\sigma)$
Applying (3.11) to (3.58) and (3.60) gives
(3.62) $\|P_{k}[(\bm{A}(s)\cdot\bm{A}(s),\ \bm{\psi}(s)\cdot\bm{\psi}(s),\ \partial_{x}\bm{A}(s))\cdot\psi_{t}(s)]\|\lesssim B_{2}\varepsilon 2^{k}2^{-2j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}^{2}.$
For $\sigma\geq\frac{1}{5}$, applying (3.12) to (3.23) and (3.61) gives
(3.63) $\|P_{k}[(\bm{A}(s)\cdot\bm{A}(s),\ \bm{\psi}(s)\cdot\bm{\psi}(s),\ \partial_{x}\bm{A}(s))\cdot\psi_{t}(s)]\|$
$\lesssim 2^{-(\sigma-1)k}2^{-2j}\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{-j}^{2}a_{k}(\sigma).$
Combining this with (3.33) yields
(3.64) $\|K(\psi_{t})\|$ $\lesssim 2^{-(\sigma-1)k}(2^{-2j}a_{-j}^{2}+\varepsilon 2^{k-j}\langle 2^{-\frac{k+j}{2}}\rangle)\langle 2^{2k}s\rangle^{-\frac{7}{2}}a_{k}(\sigma),\ \text{for}\ \sigma\geq\frac{1}{5}$
or $\lesssim B_{2}\varepsilon^{\frac{1}{2}}2^{k}(2^{-2j}a_{-j}^{2}+2^{k-j}\langle 2^{-\frac{k+j}{2}}\rangle)\langle 2^{2k}s\rangle^{-\frac{7}{2}},$
which takes the similar form as (3.53).
Repeating the argument used for $\bm{\psi}$ gives
(3.65) $\|P_{k}\psi_{t}(s)\|_{G}$ $\lesssim 2^{-(\sigma-1)k}\langle 2^{2k}s\rangle^{-4}(b_{k}(\sigma)+\varepsilon a_{k}(\sigma)),\ \text{for}\ \sigma\geq\frac{1}{5},$
or $\lesssim(1+B_{2}\varepsilon)2^{k}\langle 2^{2k}s\rangle^{-4}\varepsilon^{\frac{1}{2}}.$
By definition (1.45), (3.65) and the above imply $a_{k}(\sigma)\lesssim b_{k}(\sigma)+\varepsilon a_{k}(\sigma),\ B_{2}\lesssim(1+B_{2}\varepsilon)$, hence $a_{k}(\sigma)\lesssim b_{k}(\sigma),\ B_{2}\lesssim 1$ since $\varepsilon$ is sufficiently small. This completes the proof. ∎
Finally, we establish bilinear estimates involving $\bm{\psi}(s)$.
Proposition 3.6.
Under the bootstrap assumptions in Proposition 1.6 we have
(3.66) $\sup_{y}\Big\|\|P_{k}\bm{\psi}(s)^{y}\|_{L^{2}_{\widehat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}w\|_{L^{2}_{\widehat{x}_{m}}}\Big\|_{L^{2}_{t,x_{m}}}$
[…]
and
(3.67) $\sup_{y}\|P_{k}\bm{\psi}(s)P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}$
[…]
Moreover, we have
(3.68) $\sup_{y}\|P_{k}\bm{\psi}(s)P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w),$
(3.69) $\sup_{y}\|P_{k}D_{x}\bm{\psi}(s)P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-(\sigma-1)k}(2^{2k}s)^{-\frac{1}{4}}\langle 2^{2k}s\rangle^{-3}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w).$
Proof.
For (3.66), Let $B_{3}$ be the smallest number in $1,\infty)$ such that
(3.70) $\sup_{y}\Big\|\|P_{k}\bm{\psi}(s)^{y}\|_{L^{2}_{\widehat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}w\|_{L^{2}_{\widehat{x}_{m}}}\Big\|_{L^{2}_{t,x_{m}}}$
[…]
we use (3.51) to expand
$\sup_{y}\Big\|\|P_{k}\bm{\psi}(s)^{y}\|_{L^{2}_{\widehat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}w\|_{L^{2}_{\widehat{x}_{m}}}\Big\|_{L^{2}_{t,x_{m}}}$
$=\sup_{y}\Big\|\|P_{k}e^{s\Delta_{x}}\bm{\psi}(0)^{y}\|_{L^{2}_{\widehat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}w\|_{L^{2}_{\widehat{x}_{m}}}\Big\|_{L^{2}_{t,x_{m}}}$
$\quad+\int_{0}^{s}\Big\|\|P_{k}e^{(s-r)\Delta_{x}}K(\bm{\psi})(r)^{y}\|_{L^{2}_{\widehat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}w\|_{L^{2}_{\widehat{x}_{m}}}\Big\|_{L^{2}_{t,x_{m}}}$
$\lesssim\langle 2^{2k}s\rangle^{-N}\sup_{y}\Big\|\|P_{k}\bm{\psi}(0)^{y}\|_{L^{2}_{\widehat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}w\|_{L^{2}_{\widehat{x}_{m}}}\Big\|_{L^{2}_{t,x_{m}}}$
(3.71) $\quad+\int_{0}^{s}\langle 2^{2k}(s-r)\rangle^{-N}\sup_{y}\Big\|\|P_{k}K(\bm{\psi})(r)^{y}\|_{L^{2}_{\widehat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}w\|_{L^{2}_{\widehat{x}_{m}}}\Big\|_{L^{2}_{t,x_{m}}}.$
The first term is controlled by bootstrap assumptions and Proposition 2.5. For the second, decompose $K(\bm{\psi})$ as
(3.72) $P_{k}K(\bm{\psi})=\sum_{k_{1},k_{2}}P_{k}L(P_{k_{1}}Q,P_{k_{2}}\bm{\psi}),\ \text{where}\ Q=(2^{k}\bm{A},\partial_{x}\bm{A},\bm{A}\cdot\bm{A},\bm{\psi}\cdot\bm{\psi}).$
Using (3.24), (3.28), (3.31) and $\sum a_{k}^{2}\leq\varepsilon$, we have
(3.73) $\|P_{k_{1}}Q(s)\|_{G}\lesssim\varepsilon^{\frac{1}{2}}2^{\max\{k,k_{1}\}}2^{-\sigma k_{1}}(2^{2k_{1}}s)^{-\frac{1}{8}}\langle 2^{2k_{1}}s\rangle^{-3}b_{k_{1}}(\sigma).$
Denote by $k_{\min},k_{\text{med}}$ and $k_{\max}$ the minimum, median and maximum of the tuple $(k_{1},k_{2},k)$ respectively. For summation over $|k_{\text{med}}-k_{\max}|\leq 4$ and $|k_{2}-k|\geq 100$, we have
[…]
$\qquad\qquad\qquad\ \|P_{k_{1}}Q(r)\|_{L^{\infty}_{t,x_{m}}L^{2}_{\widehat{x}_{m}}}\sup_{y}\Big\|\|P_{k_{2}}\bm{\psi}(r)^{y}\|_{L^{2}_{\widehat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}w\|_{L^{2}_{\widehat{x}_{m}}}\Big\|_{L^{2}_{t,x_{m}}}\mathrm{d}r$
$\lesssim B_{3}2^{-\frac{k^{\prime}}{2}}\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{1},k_{2}}2^{\max\{k,k_{1}\}}2^{\frac{k_{1}+k_{\min}}{2}}2^{-\sigma\max\{k_{1},k_{2}\}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)$
$\qquad\ \int_{0}^{s}\langle 2^{2k}(s-r)\rangle^{-N}(2^{2k_{1}}r)^{-\frac{1}{8}}\langle 2^{2k_{1}}r\rangle^{-3}\langle 2^{2k_{2}}r\rangle^{-3}\mathrm{d}r$
$\lesssim B_{3}2^{-\sigma k}2^{-\frac{k^{\prime}}{2}}\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{1},k_{2}}2^{\max\{k,k_{1}\}}2^{\frac{k_{1}+k_{\min}}{2}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)$
$\qquad\langle 2^{2k}s\rangle^{-3}\int_{0}^{s}\langle 2^{2k}(s-r)\rangle^{3-N}(2^{2k_{1}}r)^{-\frac{1}{8}}\langle 2^{2\min\{k_{1},k_{2}\}}r\rangle^{-3}\mathrm{d}r$
$\lesssim B_{3}2^{-\frac{k^{\prime}}{2}}2^{-\sigma k}\langle 2^{2k}s\rangle^{-3}\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{1},k_{2}}2^{\frac{k_{\min}-k_{\max}}{2}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)$
(3.74) $\lesssim B_{3}2^{-\frac{k^{\prime}}{2}}2^{-\sigma k}\langle 2^{2k}s\rangle^{-3}\mathcal{D}(P_{k^{\prime}}w)b_{k}b_{k}(\sigma)\lesssim B_{3}\varepsilon^{\frac{1}{2}}2^{-\frac{k^{\prime}}{2}}2^{-\sigma k}\langle 2^{2k}s\rangle^{-3}\mathcal{D}(P_{k^{\prime}}w)b_{k}(\sigma),$
where we use the fact that $\max\{k_{1},k_{2}\}\geq k-4$ and $\max\{\min\{k_{1},k_{2}\},k\}\geq k_{\max}-4$.
For summmation over $|k_{2}-k|\leq 100$, we have $|k_{1}-k^{\prime}|\leq 200$ and thus
[…]
$\qquad\qquad\qquad\ \|\Xi_{m}(D)P_{k^{\prime}}w\|_{L^{\infty}_{t}L^{2}_{x}}2^{\frac{k}{2}}\|P_{k_{2}}\bm{\psi}(r)\|_{G}\|P_{k_{1}}Q(r)\|_{G}\mathrm{d}r$
$\lesssim\int_{0}^{s}\langle 2^{2k}(s-r)\rangle^{-N}2^{\frac{k}{2}+k^{\prime}}2^{-\sigma k^{\prime}}\mathcal{D}(P_{k^{\prime}}w)b_{k^{\prime}}b_{k^{\prime}}(\sigma)(2^{2k^{\prime}}r)^{-\frac{1}{8}}\langle 2^{2k^{\prime}}r\rangle^{-6}\mathrm{d}r$
$\lesssim\langle 2^{2k}s\rangle^{-3}\int_{0}^{s}\langle 2^{2k}(s-r)\rangle^{3-N}2^{\frac{k}{2}+k^{\prime}}2^{-\sigma k^{\prime}}\mathcal{D}(P_{k^{\prime}}w)b_{k^{\prime}}b_{k^{\prime}}(\sigma)(2^{2k^{\prime}}r)^{-\frac{1}{8}}\langle 2^{2k^{\prime}}r\rangle^{-3}\mathrm{d}r$
$\lesssim\langle 2^{2k}s\rangle^{-3}2^{-\frac{k^{\prime}}{2}}2^{-\sigma k^{\prime}}\mathcal{D}(P_{k^{\prime}}w)2^{\frac{k-k^{\prime}}{2}}b_{k^{\prime}}b_{k^{\prime}}(\sigma)$
(3.75) $\lesssim\langle 2^{2k}s\rangle^{-3}2^{-\frac{k^{\prime}}{2}}2^{-\sigma k^{\prime}}\mathcal{D}(P_{k^{\prime}}w)b_{k}(\sigma).$
Combining this with (3.71) and (3.74) gives $B_{3}\lesssim 1+\varepsilon^{\frac{1}{2}}B_{3}$ and hence $B_{3}\lesssim 1$. This gives the bound (3.66).
Next we prove (3.67). We again use the formula (3.51) and repeat the arguments in (3.71). It remains to control
$\int_{0}^{s}\langle 2^{2k}(s-r)\rangle^{-N}\sup_{y}\|P_{k}K(\psi)(r)P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\mathrm{d}r$
$\lesssim\|P_{k^{\prime}}w\|_{G}\sum_{k_{1},k_{2}}\int_{0}^{s}\langle 2^{2k}(s-r)\rangle^{-N}2^{\min\{k^{\prime},k_{\min}\}}\sup_{y}\|P_{k}L(P_{k_{1}}Q(r),P_{k_{2}}\bm{\psi}(r))\|_{L^{2}_{t,x}}\mathrm{d}r$
$\lesssim\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{1},k_{2}}2^{\min\{k^{\prime},k_{\min}\}}2^{\max\{k_{1},k\}}$
$\qquad\ \int_{0}^{s}\langle 2^{2k}(s-r)\rangle^{-N}(2^{2k_{1}}r)^{-\frac{1}{8}}\langle 2^{2k_{1}}r\rangle^{-3}\langle 2^{2k_{2}}r\rangle^{-4}2^{-\sigma\max\{k_{1},k_{2}\}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)\mathrm{d}r$
$\lesssim\mathcal{D}(P_{k^{\prime}}w)2^{-\sigma k}\langle 2^{2k}s\rangle^{-3}\sum_{k_{1},k_{2}}2^{\min\{k^{\prime},k_{\min}\}-k_{\max}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)$
$\lesssim\mathcal{D}(P_{k^{\prime}}w)2^{-\sigma k}2^{-\frac{k^{\prime}-k}{2}}\langle 2^{2k}s\rangle^{-3}b_{k}b_{k}(\sigma),$
where we bound the integral as in (3.74). This gives (3.67).
It follows from (3.66) and (3.67) that
(3.76) $\sup_{y}\|P_{k}\bm{\psi}(s)P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w),$
(3.77) $\sup_{y}\|P_{k}\partial_{x}\bm{\psi}(s)P_{k^{\prime}}\bm{\psi}^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-(\sigma-1)k}\langle 2^{2k}s\rangle^{-4}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}\bm{\psi}),$
which gives the bound (3.68) and (3.69) once we prove that
(3.78) $\sup_{y}\|P_{k}(\bm{A}(s)\bm{\psi}(s))P_{k^{\prime}}\bm{\psi}^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\frac{|k-k^{\prime}|}{2}}2^{-(\sigma-1)k}(2^{2k}s)^{-\frac{1}{4}}\langle 2^{2k}s\rangle^{-3}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}\bm{\psi}).$
If $k\leq k^{\prime}$, we apply Bony calculus to the product $P_{k}(\bm{A}(s)\bm{\psi}(s))$ and obtain
(3.79) $\sup_{y}\|P_{k}(\bm{A}(s)\bm{\psi}(s))P_{k^{\prime}}\bm{\psi}^{y}\|_{L^{2}_{t,x}}$
$\lesssim\sum_{k_{1},k_{2},m}\big\|\|P_{k}(P_{k_{1}}\bm{A}(s)P_{k_{2}}\bm{\psi}(s))\|_{L^{\infty}_{\hat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}\bm{\psi}^{y}\|_{L^{2}_{\hat{x}_{m}}}\big\|_{L^{2}_{t,x_{m}}}$
$\lesssim\sum_{k_{1},k_{2},m}2^{\frac{1}{2}(\min\{k,k_{1},k_{2}\}+k)}\|P_{k_{1}}\bm{A}(s)\|_{L^{\infty}_{t,x_{m}}L^{2}_{\hat{x}_{m}}}\big\|\|P_{k_{2}}\bm{\psi}(s)\|_{L^{2}_{\hat{x}_{m}}}\|\Xi_{m}(D)P_{k^{\prime}}\bm{\psi}^{y}\|_{L^{2}_{\hat{x}_{m}}}\big\|_{L^{2}_{t,x_{m}}}$
$\lesssim 2^{-\frac{k^{\prime}}{2}}\sum_{k_{1},k_{2}}2^{\frac{1}{2}(\min\{k,k_{1},k_{2}\}+k)}2^{\frac{k_{1}}{2}}2^{-\sigma\max\{k_{1},k_{2}\}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)\mathcal{D}(P_{k^{\prime}}\bm{\psi})$
$\qquad\qquad\qquad(2^{2k_{1}}s)^{-\frac{1}{8}}\langle 2^{2k_{1}}s\rangle^{-\frac{13}{4}}\langle 2^{2k_{2}}s\rangle^{-4},$
where we use the following bound for $\bm{A}(s)$ derived from (3.24) and (3.27):
(3.80) $\|P_{k_{1}}\bm{A}(s)\|_{G}\lesssim 2^{-\sigma k_{1}}(2^{2k_{1}}s)^{-\frac{1}{8}}\langle 2^{2k_{1}}s\rangle^{-\frac{13}{4}}b_{k_{1}}(\sigma).$
For the High-Low and Low-High interactions, we have
[…]
$\lesssim 2^{\frac{k-k^{\prime}}{2}}2^{-\sigma k}b_{k}(\sigma)2^{\frac{3k}{4}}s^{-\frac{1}{8}}\langle 2^{2k}s\rangle^{-\frac{13}{4}}\mathcal{D}(P_{k^{\prime}}\bm{\psi})$
$\lesssim 2^{\frac{k-k^{\prime}}{2}}2^{-(\sigma-1)k}b_{k}(\sigma)(2^{2k}s)^{-\frac{1}{4}}\langle 2^{2k}s\rangle^{-3}\mathcal{D}(P_{k^{\prime}}\bm{\psi}),$
and for High-High interactions we have
[…]
$\lesssim 2^{\frac{k-k^{\prime}}{2}}2^{-(\sigma-1)k}b_{k}(\sigma)(2^{2k}s)^{-\frac{1}{4}}\langle 2^{2k_{3}}s\rangle^{-3}\mathcal{D}(P_{k^{\prime}}\bm{\psi}).$
Thus the bound for $k^{\prime}\geq k$ is consistent with the stated bound (3.69).
If $k^{\prime}\leq k$, we have
(3.81) $\sup_{y}\|P_{k}(\bm{A}(s)\bm{\psi}(s))P_{k^{\prime}}\bm{\psi}^{y}\|_{L^{2}_{t,x}}$
$\lesssim\sum_{k_{1},k_{2}}2^{\min\{k^{\prime},k_{1},k_{2}\}}\|P_{k_{1}}\bm{A}(s)\|_{G}\|P_{k_{2}}\bm{\psi}(s)\|_{G}\|P_{k^{\prime}}\bm{\psi}\|_{G}$
$\lesssim\mathcal{D}(P_{k^{\prime}}\bm{\psi})\sum_{k_{1},k_{2}}2^{\min\{k^{\prime},k_{1},k_{2}\}}2^{-\sigma\max\{k_{1},k_{2}\}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)(2^{2k_{1}}s)^{-\frac{1}{8}}\langle 2^{2k_{1}}s\rangle^{-\frac{13}{4}}\langle 2^{2k_{2}}s\rangle^{-4}$
$\lesssim 2^{-\sigma k}2^{\frac{k^{\prime}}{2}}s^{-\frac{1}{4}}\mathcal{D}(P_{k^{\prime}}\bm{\psi})\sum_{k_{1},k_{2}}2^{\frac{1}{8}(\min\{k,k_{1},k_{2}\}-\max\{k,k_{1},k_{2}\})}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)$
$\qquad\qquad(2^{2k_{1}}s)^{\frac{1}{16}}(2^{2k_{2}}s)^{\frac{1}{16}}\langle 2^{2k_{1}}s\rangle^{-\frac{13}{4}}\langle 2^{2k_{2}}s\rangle^{-4}$
$\lesssim 2^{-(\sigma-1)k}2^{\frac{k^{\prime}-k}{2}}(2^{2k}s)^{-\frac{1}{4}}\langle 2^{2k}s\rangle^{-3}b_{k}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}\bm{\psi}),$
where we use the fact that $\max\{k_{1},k_{2}\}\geq k$. This gives the stated bound when $k^{\prime}\leq k$ and completes the proof.
4. Estimates for Ishimori equation
We begin by recalling the nonlinear terms in the Ishimori system. Denote by
$\mathcal{N}_{m}$ $=-2i\mu_{l}A_{l}\partial_{l}\psi_{m}+(A_{t}+\mu_{l}(A_{l}^{2}-i\partial_{l}A_{l}))\psi_{m}-i\mu_{l}\psi_{l}\Im(\psi_{m}\bar{\psi}_{l})$
$\quad+i\epsilon_{ij}\kappa(D_{m}\psi_{l}R_{l}R_{i}A_{j}+\psi_{l}\partial_{m}(R_{l}R_{i}A_{j}))$
$=-i\mu_{l}A_{l}\partial_{l}\psi_{m}-i\mu_{l}\partial_{l}(A_{l}\psi_{m})+i\epsilon_{ij}\kappa\ \partial_{m}(\psi_{l}R_{l}R_{i}A_{j})$
$\quad\ +(A_{t}+\mu_{l}A_{l}^{2})\psi_{m}+i\epsilon_{ij}\kappa\ \psi_{l}A_{m}R_{l}R_{i}A_{j}-i\mu_{l}\psi_{l}\Im(\psi_{m}\bar{\psi}_{l})$
(4.1) $=:\mathcal{N}_{m,1}+\mathcal{N}_{m,2}$
appearing in the Ishimori equation (1.19).
Lemma 4.1.
Under the bootstrap assumptions in Proposition 1.7, we have
(4.2) $\sup_{y}\|P_{k}(\bm{A},\bm{\psi})P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\sigma k}2^{-\frac{|k-k^{\prime}|}{2}}\mathcal{D}(P_{k^{\prime}}w)b_{k}(\sigma),$
(4.3) $\sup_{y}\|P_{k}w^{y}P_{k}[(\bm{A}\cdot\bm{A},\bm{\psi}\cdot\bm{\psi})\bm{\psi}]\|_{L^{1}_{t,x}}\lesssim\varepsilon 2^{-\sigma k}\mathcal{D}(P_{k}w)b_{k}(\sigma),$
(4.4) $\sup_{y}\|P_{k}w^{y}P_{k}(A_{t}\bm{\psi})\|_{L^{1}_{t,x}}\lesssim\varepsilon 2^{-\sigma k}\mathcal{D}(P_{k}w)b_{k}(\sigma).$
Using these bounds, we have
(4.5) $\sup_{y}\|P_{k}w^{y}P_{k}\mathcal{N}_{2,m}\|\lesssim\varepsilon 2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k}w).$
Proof.
We begin with the proof of (4.2). The bound for the $\bm{\psi}$ term
(4.6) $\|P_{k}\bm{\psi}P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\lesssim 2^{-\sigma k}2^{-\frac{|k-k^{\prime}|}{2}}\mathcal{D}(P_{k^{\prime}}w)b_{k}(\sigma)$
follows directly from the bootstrap assumptions (1.7) and the bilinear estimate (2.6). Thus it remains to estimate the product involving the connection coefficient $\bm{A}$. Using the integral representation (1.27) for $\bm{A}$, we need to control
(4.7) $\sup_{y}\|P_{k}\bm{A}P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\lesssim\sum_{k_{1},k_{2}}\int_{0}^{\infty}\|P_{k}(P_{k_{1}}\bm{\psi}(s)P_{k_{2}}[D_{x}\bm{\psi}(s)])P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\mathrm{d}s.$
We consider two cases based on the relative sizes of $k$ and $k^{\prime}$.
If $k\leq k^{\prime}$, we use the bilinear estimate (3.69) to obtain
$\sum_{k_{1},k_{2}}\int_{0}^{\infty}\|P_{k}(P_{k_{1}}\bm{\psi}(s)P_{k_{2}}[D_{x}\bm{\psi}(s)])P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\mathrm{d}s$
$\lesssim\sum_{k_{1},k_{2}}\int_{0}^{\infty}\|P_{k_{1}}\bm{\psi}(s)\|_{L^{\infty}_{t,x}}\sup_{y}\|P_{k_{2}}(D_{x}\bm{\psi}(s))P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\mathrm{d}s$
$\lesssim\sum_{k_{1},k_{2}}2^{k_{1}+k_{2}}2^{-\frac{|k_{2}-k^{\prime}|}{2}}\mathcal{D}(P_{k^{\prime}}w)$
$\quad\quad\int_{0}^{\infty}(2^{2k_{2}}s)^{-\frac{1}{4}}\langle 2^{2k_{1}}s\rangle^{-3}\langle 2^{2k_{2}}s\rangle^{-3}2^{-\sigma\max\{k_{1},k_{2}\}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)\mathrm{d}s$
$\lesssim 2^{-\sigma k}\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{1},k_{2}}2^{-\frac{|k_{1}-k_{2}|}{2}}2^{-\frac{|k_{2}-k^{\prime}|}{2}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)$
$\lesssim 2^{-\sigma k}\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{3}\leq k}2^{-\frac{|k_{3}-k|}{2}}\max\{2^{-\frac{|k-k^{\prime}|}{2}},2^{-\frac{|k_{3}-k^{\prime}|}{2}}\}b_{k_{3}}b_{k}(\sigma)$
$\quad\quad+2^{-\sigma k}\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{3}\geq k}2^{-\frac{|k_{3}-k^{\prime}|}{2}}b_{k_{3}}b_{k_{3}}(\sigma)$
(4.8) $\lesssim\varepsilon^{\frac{1}{2}}2^{-\sigma k}2^{-\frac{|k-k^{\prime}|}{2}}\mathcal{D}(P_{k^{\prime}}w)b_{k}(\sigma),$
which yields the desired bound (4.2) in this case.
We further decompose the product $P_{k}(P_{k_{1}}\bm{\psi}(s)P_{k_{2}}[D_{x}\bm{\psi}(s)])$ according to Bony’s paraproduct structure. For the Low-High interaction $k_{1}\leq k-4$, $|k_{2}-k|\leq 8$ , we have
$\sum^{|k_{2}-k|\leq 8}_{k_{1}\leq k-4}\int_{0}^{\infty}\sup_{y}\|P_{k}(P_{k_{1}}\bm{\psi}(s)P_{k_{2}}[D_{x}\bm{\psi}(s)])P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}$
$\lesssim\sum^{|k_{2}-k|\leq 8}_{k_{1}\leq k-4}\int_{0}^{\infty}\|P_{k_{1}}\bm{\psi}(s)\|_{L^{\infty}_{t,x}}\sup_{y}\|P_{k_{2}}(D_{x}\bm{\psi}(s))P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\mathrm{d}s$
$\lesssim\sum^{|k_{2}-k|\leq 8}_{k_{1}\leq k-4}2^{k_{1}+k_{2}}2^{-\frac{|k_{2}-k^{\prime}|}{2}}2^{-\sigma k_{2}}b_{k_{1}}b_{k_{2}}(\sigma)\mathcal{D}(P_{k^{\prime}}w)\int_{0}^{\infty}(2^{2k_{2}}s)^{-\frac{1}{4}}\langle 2^{2k_{1}}s\rangle^{-3}\langle 2^{2k_{2}}s\rangle^{-3}\mathrm{d}s$
$\lesssim 2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{1}\leq 4}2^{-\frac{|k_{1}-k|}{2}}b_{k_{1}}\lesssim\varepsilon^{\frac{1}{2}}2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w).$
For the High-Low and High-High interactions, we place the high-frequency term in $L^{2}$ and use the pointwise bound for $P_{k^{\prime}}w$ to obtain
$\sum_{k_{1}\geq k-4}^{k_{2}}\int_{0}^{\infty}\sup_{y}\|P_{k}(P_{k_{1}}\bm{\psi}(s)P_{k_{2}}[D_{x}\bm{\psi}(s)])P_{k^{\prime}}w^{y}\|_{L^{2}_{t,x}}\mathrm{d}s$
$\lesssim\sum_{k_{1}\geq k-4}^{k_{2}}\int_{0}^{\infty}\|P_{k^{\prime}}w\|_{L^{\infty}_{t,x}}\|P_{k_{1}}\bm{\psi}(s)P_{k_{2}}(D_{x}\bm{\psi}(s))\|_{L^{2}_{t,x}}\mathrm{d}s$
$\lesssim\sum_{k_{1}\geq k-4}^{k_{2}}2^{k^{\prime}+k_{2}}\mathcal{D}(P_{k^{\prime}}w)$
$\quad\quad\int_{0}^{\infty}(2^{2k_{2}}s)^{-\frac{1}{4}}\langle 2^{2k_{1}}s\rangle^{-3}\langle 2^{2k_{2}}s\rangle^{-3}2^{-\sigma\max\{k_{1},k_{2}\}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma)\mathrm{d}s$
$\lesssim 2^{-\sigma k}2^{-\frac{k-k^{\prime}}{2}}\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{1}\geq k-4}^{k_{2}}2^{-\frac{|k_{1}-k_{2}|}{2}}2^{-\frac{|k_{1}-k|}{2}}b_{\min\{k_{1},k_{2}\}}b_{\max\{k_{1},k_{2}\}}(\sigma),$
(4.9) $\lesssim\varepsilon^{\frac{1}{2}}2^{-\sigma k}2^{-\frac{|k-k^{\prime}|}{2}}\mathcal{D}(P_{k^{\prime}}w)b_{k}(\sigma),$
where we use the fact that $|k_{2}-\min\{k_{1},k_{2}\}|\leq 10$ in the High-Low and High-High regime. Combining both subcases gives the bound (4.2) when $k^{\prime}\leq k$.
We next prove (4.3). We focus on the representative term $P_{k}w^{y}P_{k}(\bm{A}\cdot\bm{A}\cdot\bm{\psi})$ in (4.3), as estimates for other terms are similar. Applying Bony’s paraproduct decomposition, we obtain
$\sup_{y}\|P_{k}w^{y}P_{k}(\bm{A}\cdot\bm{A}\cdot\bm{\psi})\|_{L^{1}_{t,x}}\lesssim\sum_{k_{i}}\sup_{y}\|P_{k}w^{y}P_{k}(P_{k_{1}}\bm{A}\cdot P_{k_{2}}\bm{A}\cdot P_{k_{3}}\bm{\psi})\|_{L^{1}_{t,x}}$
$\lesssim\sum_{k_{i}}\min\{2^{-\frac{1}{2}(|k-k_{1}|+|k_{2}-k_{3}|)},2^{-\frac{1}{2}(|k-k_{2}|+|k_{1}-k_{3}|)}\}2^{-\sigma k_{\max}}b_{k_{\min}}b_{k_{\text{med}}}b_{k_{\max}}(\sigma)\mathcal{D}(P_{k}w)$
$\lesssim 2^{-\sigma k}\mathcal{D}(P_{k}w)\sum_{k_{i}}2^{-\frac{1}{8}(|k-k_{1}|+|k_{1}-k_{2}|+|k_{2}-k_{3}|)}b_{k_{\min}}b_{k_{\text{med}}}b_{k_{\max}}(\sigma)$
(4.10) $\lesssim\varepsilon 2^{-\sigma k}\mathcal{D}(P_{k}w)b_{k}(\sigma),$
where $(k_{\min},k_{\text{med}},k_{\max})$ is the increasing rearrangement of $(k_{1},k_{2},k_{3})$ and the off-diagonal factor comes from applying (4.2) to the pairs $(P_{k}w^{y},P_{k_{1}}\bm{A})$ and $(P_{k_{2}}\bm{A},P_{k_{3}}\bm{\psi})$ in the two possible bilinear groupings.
Finally, we prove (4.4). Using the identity $A_{t}=-\int_{s}^{\infty}\Im(\overline{\psi_{t}}D_{l}\psi_{l})(r)\mathrm{d}r$, we need to distinguish between two regularity regimes for $\psi_{t}$, as given by (3.49).
For $\sigma\leq\frac{1}{5}$, we shall use the bound $\|P_{k}\psi_{t}(s)\|\lesssim\varepsilon^{\frac{1}{2}}2^{k}\langle 2^{2k}s\rangle^{-4}$ to obtain
$\|P_{k}w^{y}P_{k}(A_{t}\bm{\psi})\|_{L^{1}_{t,x}}\lesssim\sum_{k_{i}}\int_{0}^{+\infty}\|P_{k}w^{y}P_{k}(P_{k_{1}}\psi_{t}(s)P_{k_{2}}(D_{x}\bm{\psi}(s))P_{k_{3}}\bm{\psi})\|_{L^{1}_{t,x}}\mathrm{d}s$
$\lesssim\sum_{k_{i}}2^{k_{1}+k_{2}}\min\{2^{-\frac{|k_{2}-k_{3}|}{2}},2^{-\frac{|k-k_{2}|}{2}}\}2^{-\sigma k_{3}}\varepsilon^{\frac{1}{2}}b_{k_{2}}b_{k_{3}}(\sigma)\mathcal{D}(P_{k}w)$
$\qquad\qquad\int_{0}^{\infty}(2^{2k_{2}}s)^{-\frac{1}{4}}\langle 2^{2k_{1}}s\rangle^{-3}\langle 2^{2k_{2}}s\rangle^{-3}\mathrm{d}s$
$\lesssim\varepsilon^{\frac{1}{2}}2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k}w)\sum_{k_{i}}2^{-\frac{|k_{1}-k_{2}|}{2}}2^{(\sigma+\delta)|k-k_{3}|}\min\{2^{-\frac{|k_{2}-k_{3}|}{2}},2^{-\frac{|k-k_{2}|}{2}}\}b_{k_{2}}$
(4.11) $\lesssim\varepsilon 2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k}w),$
where the factor $\min\{2^{-\frac{|k_{2}-k_{3}|}{2}},2^{-\frac{|k-k_{2}|}{2}}\}$ comes from applying (3.69) either to the pair $(P_{k}w^{y},P_{k_{2}}(D_{x}\bm{\psi}))$ or to the pair $(P_{k_{3}}\bm{\psi},P_{k_{2}}(D_{x}\bm{\psi}))$, whichever yields better decay.
For $\sigma\geq\frac{1}{5}$, we shall use the bound $\|P_{k}\psi_{t}(s)\|\lesssim 2^{-(\sigma-1)k}\langle 2^{2k}s\rangle^{-4}b_{k}(\sigma)$ to obtain
$\|P_{k}w^{y}P_{k}(A_{t}\bm{\psi})\|_{L^{1}_{t,x}}\lesssim\sum_{k_{i}}\int_{0}^{+\infty}\|P_{k}w^{y}P_{k}(P_{k_{1}}\psi_{t}(s)P_{k_{2}}(D_{x}\bm{\psi}(s))P_{k_{3}}\bm{\psi})\|_{L^{1}_{t,x}}\mathrm{d}s$
$\lesssim\sum_{k_{i}}2^{k_{1}+k_{2}}\min\{2^{-\frac{|k_{2}-k_{3}|}{2}},2^{-\frac{|k-k_{2}|}{2}}\}2^{-\sigma k_{\max}}b_{k_{\min}}b_{k_{\text{med}}}b_{k_{\max}}(\sigma)\mathcal{D}(P_{k}w)$
$\qquad\qquad\int_{0}^{\infty}(2^{2k_{2}}s)^{-\frac{1}{4}}\langle 2^{2k_{1}}s\rangle^{-3}\langle 2^{2k_{2}}s\rangle^{-3}\mathrm{d}s$
$\lesssim 2^{-\sigma k}\mathcal{D}(P_{k}w)\sum_{k_{i}}2^{-|k_{1}-k_{2}|}\min\{2^{-\frac{|k_{2}-k_{3}|}{2}},2^{-\frac{|k-k_{2}|}{2}}\}b_{k_{\min}}b_{k_{\text{med}}}b_{k_{\max}}(\sigma)$
(4.12) $\lesssim\varepsilon 2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k}w).$
This completes the proof of (4.4).
IThe estimate (4.5) for $\mathcal{N}_{2,m}$ follows directly from (4.4) and (4.3), combined with the fact that the remaining term $P_{k}w^{y}P_{k}(\psi_{l}A_{m}R_{l}R_{i}A_{j})$ can be controlled in a similar manner to (4.3), since
(4.13) $P_{k}w^{y}P_{k}(\psi_{l}A_{m}R_{l}R_{i}A_{j})=\sum_{k_{i}}P_{k}w^{y}L(P_{k_{1}}\bm{\psi},P_{k_{2}}\bm{A},P_{k_{3}}A_{j})$
and (4.2) applies well to the $L$-notation product terms. This completes the proof. ∎
Lemma 4.2.
Under the bootstrap assumptions, the derivative part of the nonlinearity satisfies
(4.14) $\sup_{y}\|P_{k}w^{y}P_{k}\mathcal{N}_{1,m}\|\lesssim\varepsilon 2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k}w)$
Proof.
We apply Bony’s paraproduct decomposition to the terms in $\mathcal{N}_{1,m}$
$\sup_{y}\|P_{k}w^{y}P_{k}\mathcal{N}_{1,m}\|$
$\lesssim\sum_{k^{\prime},k_{3}}\int_{0}^{\infty}\sup_{y,y^{\prime}}\|P_{k}w^{y}\cdot P_{k}(\partial_{x}L(P_{k^{\prime}}AP_{k_{3}}\bm{\psi}),L(P_{k^{\prime}}A\partial_{x}P_{k_{3}}\bm{\psi}))\|_{L^{1}_{t,x}}\mathrm{d}s$
$\lesssim\sum_{k_{i}}2^{\max\{k,k_{3}\}}\int_{0}^{\infty}\sup_{y,y^{\prime}}\|P_{k}w^{y}P_{k_{1}}\bm{\psi}(s)P_{k_{2}}D_{x}\bm{\psi}P_{k_{3}}\bm{\psi}^{y^{\prime}}\|_{L^{1}_{t,x}}$
$\lesssim\sum_{k_{i}}2^{\max\{k,k_{3}\}+k_{2}}\min\{2^{-\frac{1}{2}(|k-k_{1}|+|k_{2}-k_{3}|)},2^{-\frac{1}{2}(|k-k_{2}|+|k_{1}-k_{3}|)}\}$
$\quad\cdot\int_{0}^{\infty}(2^{2k_{2}})^{-\frac{1}{4}}\langle 2^{2k_{1}}s\rangle^{-3}\langle 2^{2k_{2}}s\rangle^{-3}\mathrm{d}s\cdot 2^{-\sigma k_{\max}}b_{k_{\min}}b_{k_{\text{med}}}b_{k_{\max}}(\sigma)\mathcal{D}(P_{k}w)$
$\lesssim\sum_{k_{i}}2^{\max\{k,k_{3}\}+k_{2}-2\max\{k_{1},k_{2}\}}\min\{2^{-\frac{1}{2}(|k-k_{1}|+|k_{2}-k_{3}|)},2^{-\frac{1}{2}(|k-k_{2}|+|k_{1}-k_{3}|)}\}$
$\quad\ b_{k_{\min}}b_{k_{\text{med}}}b_{k_{\max}}(\sigma)\mathcal{D}(P_{k}w)$
(4.15) $\lesssim 2^{-\sigma k}\mathcal{D}(P_{k}w)\sum_{k_{i}}C(k,k_{1},k_{2},k_{3})b_{k_{\min}}b_{k_{\text{med}}}b_{k_{\max}}(\sigma),$
where $(k_{\min},k_{\text{med}},k_{\max})$ is the increasing rearrangement of $(k_{1},k_{2},k_{3})$, and the factor $C(k,k_{1},k_{2},k_{3})$ is defined as
(4.16) $C(k,k_{1},k_{2},k_{3})=1_{\{|k_{\max}-k|\leq 10\}\cup\{k_{\max}\geq k+10,\ |k_{\max}-k_{\text{med}}|\leq 10\}}$
$2^{\max\{k,k_{3}\}-\max\{k_{1},k_{2}\}}\min\{2^{-\frac{1}{2}(|k-k_{1}|+|k_{2}-k_{3}|)},2^{-\frac{1}{2}(|k-k_{2}|+|k_{1}-k_{3}|)}\}.$
The indicator function restricts to the two main frequency interaction scenarios: either the highest frequency is close to $k$, or it is significantly larger than $k$ but close to the second highest.
To bound the sum in (4.15), we consider two subcases based on the relationship between $\max\{k,k_{3}\}$ and $\max\{k_{1},k_{2}\}$.
If $\max\{k,k_{3}\}\leq\max\{k_{1},k_{2}\}+10$, then
$C(k,k_{1},k_{2},k_{3})\lesssim 2^{-\frac{|\Delta k|}{4}},\ \Delta k:=\max\{k,k_{1},k_{2},k_{3}\}-\min\{k,k_{1},k_{2},k_{3}\},$
which provides strong off-diagonal decay.
If $\max\{k,k_{3}\}\geq\max\{k_{1},k_{2}\}+10$, then the factor vanishes unless $k_{\max}=k_{3}$ and $|k_{3}-k|\leq 10$. In this scenario we have
(4.17) $C(k,k_{1},k_{2},k_{3})\lesssim 2^{k-\max\{k_{1},k_{2}\}}\cdot 2^{-\frac{1}{2}(|k-k_{1}|+|k-k_{2}|)}\lesssim 2^{-\frac{1}{2}|k_{1}-k_{2}|}.$
The gain comes from the mismatch between $k$ and $\max\{k_{1},k_{2}\}$.
Combining both subcases, we estimate
(4.15) $\lesssim 2^{-\sigma k}\mathcal{D}(P_{k}w)\sum_{\max\{k,k_{3}\}\leq\max\{k_{1},k_{2}\}}2^{-\frac{|\Delta k|}{4}}b_{k_{\min}}b_{k_{\text{med}}}b_{k_{\max}}(\sigma)$
$\quad+2^{-\sigma k}\mathcal{D}(P_{k}w)\sum_{k_{1},k_{2}}2^{-\frac{1}{2}|k_{1}-k_{2}|}b_{k_{1}}b_{k_{2}}b_{k}(\sigma)$
(4.18) $\lesssim\varepsilon 2^{-\sigma k}\mathcal{D}(P_{k}w)b_{k}(\sigma),$
which gives (4.5) and thus completes the proof. ∎
Proposition 4.3.
Under bootstrap assumptions in Proposition 1.7, we have
(4.19) $b_{k}(\sigma)+\mathcal{D}(P_{k}\bm{\psi})\lesssim c_{k}(\sigma).$
Proof.
By the definition (1.46) of $b_{k}(\sigma)$, it suffices to prove the following two estimates:
(4.20) $\mathcal{D}(P_{k}\bm{\psi})\lesssim 2^{-\sigma k}c_{k}(\sigma),$
(4.21) $\|P_{k}\psi_{t}\|_{G}\lesssim 2^{-(\sigma-1)k}c_{k}(\sigma),\ \sigma\in[\frac{1}{5},\sigma_{1}].$
We apply the Strichartz estimate (2.11) for the ultrahyperbolic Schrödinger equation to the differentiated Ishimori equation (1.19). This yields
(4.22) $\mathcal{D}(P_{k}\bm{\psi})\lesssim 2^{-\sigma k}c_{k}(\sigma)+\sum_{m=1,2}\|P_{k}\mathcal{N}_{m}\|_{DV^{2}_{\text{UH}}}+I(P_{k}\bm{\psi},P_{k}\mathcal{N}_{m})^{\frac{1}{2}}.$
The integral $I(P_{k}\bm{\psi},\mathcal{N}_{m})^{\frac{1}{2}}$ is controlled directly by (4.5) and (4.14). Using the bootstrap assumption $\mathcal{D}(P_{k}\bm{\psi})\lesssim 2^{-\sigma k}b_{k}(\sigma)$ and $b_{k}(\sigma)\leq\varepsilon^{-\frac{1}{4}}c_{k}(\sigma)$, we obtain
(4.23) $I(P_{k}\bm{\psi},P_{k}\mathcal{N}_{m})^{\frac{1}{2}}\lesssim\mathcal{D}(P_{k}\bm{\psi})^{\frac{1}{2}}(\varepsilon 2^{-\sigma k}b_{k}(\sigma))^{\frac{1}{2}}\lesssim\varepsilon^{\frac{1}{2}}2^{-\sigma k}b_{k}(\sigma)\lesssim 2^{-\sigma k}c_{k}(\sigma).$
The norm $\|P_{k}\mathcal{N}_{m}\|_{DV^{2}_{\text{UH}}}$ is estimated by duality. For any $U^{2}_{\text{UH}}$ atom $w$ (which implies $\mathcal{D}(P_{k}w)\lesssim 1$), we have
(4.24) $\big|\int\bar{w}P_{k}\ \mathcal{N}_{m}\mathrm{d}t\mathrm{d}x\big|\lesssim\|P_{\sim k}wP_{k}\mathcal{N}_{m}\|_{L^{1}_{t,x}}\lesssim\varepsilon 2^{-\sigma k}b_{k}(\sigma)\lesssim 2^{-\sigma k}c_{k}(\sigma).$
It remains to prove the bound for $\psi_{t}$. Usinng the equation (1.17), we have
(4.25) $\psi_{t}=\partial_{x}\bm{\psi}+\bm{A}\cdot\bm{\psi}+(R^{2}\bm{A})\cdot\bm{\psi},$
where $R$ is the Riesz operators. The term $\partial_{x}\bm{\psi}$ is already controlled by the estimate for $\bm{\psi}$. The remaining terms involve products of the connection coefficient $\bm{A}$ and the field $\bm{\psi}$. To estimate these products, we use the following bound for $\bm{A}$, obtained by taking the limit $s\to 0$ in the heat flow estimate (3.42) and using the previously established fact that $a_{k}(\sigma)\lesssim b_{k}(\sigma)\lesssim\varepsilon^{\frac{1}{2}}$:
(4.26) $\|P_{k}\bm{A}\|_{G}\lesssim\varepsilon^{\frac{1}{2}}2^{-\sigma k}b_{k}(\sigma)\lesssim\varepsilon^{\frac{1}{2}}2^{-\sigma k}b_{k}(\sigma).$
Applying Lemma 3.1 to the product together with the bound (4.26) and the bootstrap bound for $\bm{\psi}$, we obtain
(4.27) $\|P_{k}(\bm{A}\cdot\bm{\psi},(R^{2}\bm{A})\cdot\bm{\psi})\|_{G}\lesssim 2^{-(\sigma-1)k}c_{k}(\sigma),\ \sigma\in[\frac{1}{5},\sigma_{1}],$
which completes the proof. ∎
Acknowledgements
Y. Zhou was partially supported by NSFC (No. 12171097).